### v1

# Geometric and topological rigidity of pinched submanifolds II

Abstract
We continue the study of the geometry and topology of compact submanifolds of arbitrary codimension in space forms satisfying a certain pinching condition involving the length of the second fundamental form and the mean curvature. Our primary focus is on four-dimensional submanifolds, where, to our surprise, both the results obtained and the methods employed differ significantly and are notably more intricate compared to those in higher dimensions. This study heavily relies on concepts from four-dimensional geometry, the geometry of Riemannian manifolds with nonnegative isotropic curvature, and the Bochner technique, each playing a crucial role. The results are sharp and extend previous results by several authors, without imposing any further assumption on either the mean curvature or the fundamental group of the submanifold.
footnotetext: 2020 Mathematics Subject Classification. 53C40, 53C42.footnotetext: Keywords and phrases. Homology groups, pinching, length of the second fundamental form, sectional, isotropic and mean curvature, Bochner-Weitzenböck operator.
1 Introduction
A fundamental problem in differential geometry is understanding the interplay between the geometry and topology of Riemannian manifolds. In the context of submanifold theory, it is particularly intriguing to investigate how the geometry or topology of submanifolds in space forms is influenced by pinching conditions involving either intrinsic or extrinsic curvature invariants.
For minimal submanifolds of spheres with a sufficiently pinched second fundamental form, an important result was first established by Simons in his seminal paper [22]. Subsequently, Chern, do Carmo, and Kobayashi [4] proved a celebrated rigidity theorem. Their work has since inspired numerous significant results in the study of pinching conditions. Notable examples include [1, 9, 12, 17, 21, 23, 28].
In our previous paper [24], we investigated the geometric and topological rigidity of submanifolds $f\colon M^{n}\to\mathbb{Q}_{c}^{n+m}$ satisfying the pinching condition
$S\leq a(n,k,H,c)$ ($\ast$)
pointwise. Here, $k$ is an integer with $21\leq k\leq n/$ 2, and the function $a$ is defined as
$a(n,k,t,c)=nc+\frac{n^{3}t^{2}}{2k(n-k)}-\frac{n|n-2k|}{2k(n-k)}t\sqrt{n^{2}t^{2}+4ck(n-k)},\;t,c\geq 0.$
Furthermore, $S$ denotes the squared length of the second fundamental form $\alpha_{f}\colon TM\times TM\to N_{f}M$, which takes values in the normal bundle $N_{f}M$, while the mean curvature is defined as the length $H=\|\mathcal{H}\|$ of the mean curvature vector field given by $\mathcal{H}=(\mathrm{tr}\,\alpha_{f})/n$, where $\mathrm{tr}$ denotes the trace. The ambient space $\mathbb{Q}_{c}^{n+m}$ is the $(n+m)$-dimensional, complete, simply connected space form of constant curvature $c$. For simplicity, we assume that $c\in\{0,1\}$, unless stated otherwise. Thus $\mathbb{Q}_{c}^{n+m}$ is the Euclidean space $\mathbb{R}^{n+m}$ ($c=0$), or the unit sphere $\mathbb{S}^{n+m}$ ($c=1$).
A direct computation shows that the standard embedding of a torus
$\mathbb{T}^{n}_{k}(r)=\mathbb{S}^{k}(r)\times\mathbb{S}^{n-k}(\sqrt{1-r^{2}}),1\leq k\leq n-1,$
into the unit sphere $\mathbb{S}^{n+1}$, where $\mathbb{S}^{k}(r)$ denotes the $k$-dimensional sphere of radius $r<1$, satisfies ($\ast$1) as an equality if $r\geq\sqrt{k/n}$, or $n=2k$. Otherwise, $S>a(n,k,H,1)$.
The pinching condition ($\ast$1) has been primarily studied for the lowest allowed $k$ (see for instance, [25, 27, 28, 30]). Shiohama and Xu [21] proved that compact submanifolds in space forms of nonnegative curvature are homeomorphic to a sphere, provided that ($\ast$1) holds as a strict inequality at any point for $k=1$. In [23], a homology vanishing result was established for submanifolds in space forms of nonnegative curvature that satisfy the strict inequality in ($\ast$1) pointwise for an integer $11\leq k\leq n-$ 1. The approach in that work relied on the nonexistence of stable currents, as shown by Lawson and Simons [10], under specific upper bounds of the second fundamental form.
In our recent paper [24], we investigated the geometric and topological rigidity of submanifolds of dimension $n\geq 5$ that satisfy the pinching condition ($\ast$1), without requiring the condition to be strict or imposing any assumptions on the mean curvature. We proved the following result, which shows that the pinching condition either enforces the vanishing of homology in a range of intermediate dimensions or uniquely determines the pinched submanifold up to congruence.
Theorem 1.
Let $f\colon M^{n}\to\mathbb{Q}_{c}^{n+m}$, $n\geq 5,c\geq 0$, be a substantial isometric immersion of a compact oriented Riemannian manifold. Assume that the inequality (italic-$\ast$1) is satisfied for an integer $22\leq k\leq n/$ 2 at any point. Then, either the homology groups of $M^{n}$ satisfy
$H_{p}(M^{n};\mathbb{Z})=0\;\,\text{for all}\;\,k\leq p\leq n-k\;\,\text{and}\;\,H_{k-1}(M^{n};\mathbb{Z})\cong\mathbb{Z}^{\beta_{k-1}(M)},$
where $\beta_{k-1}(M)$ is the $(k-1)$-th Betti number, or equality holds in (italic-$\ast$1) at any point and one of the following assertions holds:
$(i)$ $M^{n}$ is isometric to a Clifford torus $\mathbb{T}^{n}_{p}(\sqrt{p/n}),k\leq p\leq n/2$, and $f$ is the standard minimal embedding into $\mathbb{S}^{n+1}$.
$(ii)$ $M^{n}$ is isometric to a torus $\mathbb{T}^{n}_{k}(r)$ with $r>\sqrt{k/n}$ and $f$ is the standard embedding into $\mathbb{S}^{n+1}$.
$(iii)$ $M^{n}$ is isometric to a torus $\mathbb{S}^{k}(r)\times\mathbb{S}^{k}(\sqrt{R^{2}-r^{2}})$ and $f$ is a composition $f=j\circ g$, where $g\colon M^{n}\to\mathbb{S}^{n+1}(R)$ is the standard embedding of the torus $\mathbb{S}^{k}(r)\times\mathbb{S}^{k}(\sqrt{R^{2}-r^{2}})$ into a sphere $\mathbb{S}^{n+1}(R)$ and $j\colon\mathbb{S}^{n+1}(R)\to\mathbb{Q}_{c}^{n+2}$ is an umbilical inclusion with $c=0,1$ and $R<1$ if $c=1$.
This paper focuses on investigating the case of four-dimensional submanifolds, which was not addressed in Theorem 1. As will become clear, both the results obtained and the methods utilized differ substantially and are significantly more intricate than those used in the proof of Theorem 1 for higher dimensions. This study heavily relies on concepts from four-dimensional geometry, the geometry of Riemannian manifolds with nonnegative isotropic curvature, and the Bochner technique, each playing an essential role. A key observation is that four-dimensional submanifolds satisfying the pinching condition ($\ast$1) possess nonnegative isotropic curvature, a concept first introduced by Micallef and Moore [14].
Throughout the paper, all submanifolds under consideration are assumed to be connected. We recall that a submanifold is called substantial if it is not contained in a proper totally geodesic submanifold of the ambient space. The main result of the paper can be stated as follows.
Theorem 2.
Let $f\colon M^{4}\to\mathbb{Q}_{c}^{4+m},c\geq 0$, be a substantial isometric immersion of a compact, oriented Riemannian four-manifold. Suppose the inequality (italic-$\ast$1) is satisfied for $k=2$ at every point. Then, one of the following assertions holds:
$(i)$ $M^{4}$ is diffeomorphic to $\mathbb{S}^{4}$, or
$(ii)$ The universal cover of $M^{4}$ is isometric to a Riemannian product ×$\mathbb{R}\times N$, where $N$ is diffeomorphic to $\mathbb{S}^{3}$ with nonnegative Ricci curvature, or
$(iii)$ Equality holds in (italic-$\ast$1) for $k=2$ at every point, and one of the following holds:
  1. (a)
    $M^{4}$ is isometric to a torus $\mathbb{S}^{2}(r)\times\mathbb{S}^{2}(\sqrt{R^{2}-r^{2}})$ and $f$ is a composition $f=j\circ g$, where $g\colon M^{4}\to\mathbb{S}^{5}(R)$ is the standard embedding of the torus $\mathbb{S}^{2}(r)\times\mathbb{S}^{2}(\sqrt{R^{2}-r^{2}})$ into a sphere $\mathbb{S}^{5}(R)$, and $j\colon\mathbb{S}^{5}(R)\to\mathbb{Q}_{c}^{6}$ is an umbilical inclusion with $R\leq 1$ if $c=1$.
  2. (b)
    $M^{4}$ is isometric to the projective plane $\mathord{\mathbb{C}}\mathord{\mathbb{P}}_{r}^{2}$ of constant holomorphic curvature $4/3r^{2}$ with $r=1/\sqrt{c+H^{2}}$ and $f=j\circ g$, where $g$ is the standard embedding of $\mathord{\mathbb{C}}\mathord{\mathbb{P}}_{r}^{2}$ into $\mathbb{S}^{7}(r)$, and $j\colon\mathbb{S}^{7}(r)\to\mathbb{Q}_{c}^{8}$ is an umbilical inclusion.
The standard isometric embedding of the torus $\mathbb{S}^{1}(r)\times\mathbb{S}^{3}(\sqrt{1-r^{2}})$ into $\mathbb{S}^{5}$ clearly satisfies ($\ast$1) for $k=2$, provided $r\geq 1/2$, and its fundamental group is infinite. In the final section, we present a method for constructing geometrically distinct isometric immersions of manifolds that are diffeomorphic to either $\mathbb{S}^{4}$ or the torus $\mathbb{S}^{1}\times\mathbb{S}^{3}$, while also satisfying ($\ast$1) for $k=2$. These examples demonstrate that Theorem 2 is indeed sharp.
For submanifolds of arbitrary dimension that satisfy ($\ast$1) for $k=2$, we prove the following result.
Corollary 3.
Let $f\colon M^{n}\to\mathbb{Q}_{c}^{n+m}$, $n\geq 4,c\geq 0$, be a substantial isometric immersion of a compact, oriented Riemannian manifold. Assume that the inequality (italic-$\ast$1) is satisfied for $k=2$ at every point. Then, one of the following holds:
$(i)$ $M^{4}$ is diffeomorphic to $\mathbb{S}^{4}$, or
$(ii)$ The universal cover of $M^{4}$ is isometric to a Riemannian product ×$\mathbb{R}\times N$, where $N$ is diffeomorphic to $\mathbb{S}^{3}$ with nonnegative Ricci curvature, or
$(iii)$ If $n\geq 5$, the homology of $M^{n}$ is given by
$H_{p}(M^{n};\mathbb{Z})=0\;\,\text{for}\;\,2\leq p\leq n-2\;\,\text{and}\;\,H_{p}(M^{n};\mathbb{Z})\cong\mathbb{Z}^{\beta_{1}(M)}\;\,\text{for}\;\,p=1,n-1,$
or equality holds in (italic-$\ast$1) for $k=2$ at every point, and the submanifold is as described in parts $(i)-(iii)$ of Theorem 1, or as in part $(iii)$ of Theorem 2.
The results presented here extend the findings of previous studies [1, 2, 12, 28, 29, 30], without requiring additional assumptions on either the mean curvature or the fundamental group of the submanifold. Furthermore, the examples provided in the final section demonstrate the sharpness of Corollary 3.
Submanifolds of arbitrary dimension satisfying the pinching condition ($\ast$1) for $k=1$ were studied in [24]. Together with the above results, this provides a comprehensive understanding of submanifolds satisfying ($\ast$1) for the two smallest admissible values of the integer $k$.
2 Background
2.1 Geometry of 4-dimensional manifolds
In this section, we collect basic facts about four-dimensional geometry. For a detailed exposition of the subject, we refer the reader to [11, 20].
Let $(M,{\langle}\cdot,\cdot{\rangle})$ be an oriented Riemannian manifold of dimension $n=4$ with Levi-Civita connection $\nabla$ and curvature tensor $R$ given by
$R(X,Y)=[\nabla_{X},\nabla_{Y}]-\nabla_{[X,Y]},\;X,Y\in\mathcal{X}(M).$
The Ricci tensor of $(M,{\langle}\cdot,\cdot{\rangle})$ is defined by
$\mbox{Ric}(X,Y)=\sum_{i}{\langle}R(X,E_{i})E_{i},Y{\rangle},\;X,Y\in\mathcal{X}(M),$
where $\{E_{i}\}_{1\leq i\leq 4}$ is a local orthonormal frame.
At any point $x\in M$, we consider the Bochner-Weitzenböck operator $\mathcal{B}^{[2]}$ as an endomorphism of the space of 2-vectors $\Lambda^{2}T_{x}M$ at $x$ given by
${\langle}{\langle}\mathcal{B}^{[2]}(v_{1}\wedge v_{2}),w_{1}\wedge w_{2}{\rangle}{\rangle}=$ $\;\mbox{Ric}(v_{1},w_{1}){\langle}v_{2},w_{2}{\rangle}+\mbox{Ric}(v_{2},w_{2}){\langle}v_{1},w_{1}{\rangle}$
$-\mbox{Ric}(v_{1},w_{2}){\langle}v_{2},w_{1}{\rangle}-\mbox{Ric}(v_{2},w_{1}){\langle}v_{1},w_{2}{\rangle}$
$-2{\langle}R(v_{1},v_{2})w_{2},w_{1}{\rangle},$ (1)
[…]
${\langle}{\langle}v_{1}\wedge v_{2},w_{1}\wedge w_{2}{\rangle}{\rangle}=\det({\langle}v_{i},w_{j}{\rangle}).$
The Bochner-Weitzenböck operator is a self-adjoint operator. If $X\in\Lambda^{2}T_{x}M$, the dual 2-form $\omega$ is defined by $\omega(v,w)={\langle}{\langle}X,v\wedge w{\rangle}{\rangle}$, and we may consider $X$ as the skew-symmetric endomorphism of the tangent space at $x$ by ${\langle}X(u),v{\rangle}={\langle}{\langle}X,v\wedge w{\rangle}{\rangle}$.
[…]
$\mathcal{B}^{[2]}(\omega)(X_{1},X_{2})=\omega(\mbox{Ric}(X_{1}),X_{2})+\omega(X_{1},\mbox{Ric}(X_{2}))-\sum_{i}\omega(R(X_{1},X_{2})E_{i},E_{i}).$
Then the Bochner-Weitzenböck operator acts on $\Lambda^{2}TM$ by $\omega(\mathcal{B}^{[2]}(X_{1}\wedge X_{2}))=\mathcal{B}^{[2]}(\omega)(X_{1},X_{2})$. Taking $\omega$ to be the dual form to $w_{1}\wedge w_{2}$ yields (2.1). Throughout the paper, we will mostly identify 2-forms with 2-vectors.
We recall the Bochner-Weitzenböck formula which can be written as
${\langle}\Delta\omega,\omega{\rangle}=\frac{1}{2}\Delta\|\omega\|^{2}+\|\nabla\omega\|^{2}+{\langle}\mathcal{B}^{[2]}(\omega),\omega{\rangle}$
for any $\omega\in\Omega^{2}(M)$. This implies that any harmonic 2-form on a compact manifold is parallel provided that the Bochner-Weitzenböck operator is nonnegative.
The bundle of 2-forms of any oriented 4-dimensional Riemannian manifold $M$ decomposes as a direct sum
$\Omega^{2}(M)=\Omega_{+}^{2}(M)\oplus\Omega_{-}^{2}(M)$
of the eigenspaces of the Hodge star operator $\ast\colon\Omega^{2}(M)\to\Omega^{2}(M)$. The sections of $\Omega_{+}^{2}(M)$ are called self-dual 2-forms, whereas the ones of $\Omega_{-}^{2}(M)$ are called anti-self-dual 2-forms. Accordingly, we have the splitting
$\Lambda^{2}T_{x}M=\Lambda_{+}^{2}T_{x}M\oplus\Lambda_{-}^{2}T_{x}M$
at any point $x$, where $\Lambda_{\pm}^{2}T_{x}M$ are the eigenspaces of the Hodge star operator $\ast\colon\Lambda^{2}T_{x}M\to\Lambda^{2}T_{x}M$. Both spaces $\Lambda_{+}^{2}T_{x}M$ and $\Lambda_{-}^{2}T_{x}M$ are $\mathcal{B}^{[2]}$-invariant (see [19, Prop. 1]). Then we have accordingly the decomposition $\mathcal{B}^{[2]}=\mathcal{B}_{+}^{[2]}\oplus\mathcal{B}_{-}^{[2]}$.
Suppose now that $M$ is a compact oriented Riemannian four-manifold. The Hodge theorem guarantees that every de Rham cohomology class on $M$ has a unique harmonic representative. In particular, the space $\mathscr{H}^{2}(M)$ of harmonic 2-forms decomposes as
$\mathscr{H}^{2}(M)=\mathscr{H}^{2}_{+}(M)\oplus\mathscr{H}^{2}_{-}(M),$
where +$\mathscr{H}^{2}_{+}(M)$ and $\mathscr{H}^{2}_{-}(M)$ are the spaces of self-dual and anti-self-dual harmonic 2-forms, respectively. The dimensions of these subspaces, denoted by $\beta_{\pm}(M)=\dim\mathscr{H}^{2}_{\pm}(M)$, are oriented homotopy invariants of $M$. Their difference $\sigma=\beta_{+}(M)-\beta_{-}(M)$ is the signature of $M$, while their sum equals the second Betti number $\beta_{2}(M)$ of the manifold $M$.
2.2 Isotropic curvature
Let $(M,g)$ be a Riemannian manifold of dimension $n\geq 4$. We say that $(M,g)$ has nonnegative isotropic curvature at a point $x\in M$ if
$R_{1331}+R_{1441}+R_{2332}+R_{2442}-2R_{1234}\geq 0,$
for all orthonormal four-frames $\{e_{1},e_{2},e_{3},e_{4}\}\subset T_{x}M$. Here, we denote
[…]
where $R$ is the curvature tensor. If the strict inequality holds, we say that $(M,g)$ has positive isotropic curvature at $x$. The manifold $(M,g)$ is said to have nonnegative (or positive) isotropic curvature if it satisfies the corresponding condition at every point and for all orthonormal four-frames.
The following result is well known (see for instance [16]).
Lemma 4.
For any four-dimensional Riemannian manifold $M$, the non-negativity of the isotropic curvature at a point $x\in M$ is equivalent to the non-negativity of the Bochner-Weitzenböck operator $\mathcal{B}^{[2]}$ at $x$.
2.3 The pinching condition and the Lawson-Simons inequality
Lawson and Simons [10] proved that specific bounds on the second fundamental form for submanifolds of spheres can result to the vanishing of homology groups with integer coefficients. By employing the second variation of area, they effectively ruled out stable minimal currents in certain dimensions. Consequently, since area minimization can be performed within a homology class, this result implies the trivialization of integral homology.
[…]
Theorem 5.
([6, 10, 26]) Let $f\colon M^{n}\to\mathbb{Q}_{c}^{n+m},n\geq 4,c\geq 0$, be an isometric immersion of a compact Riemannian manifold and $p$ be an integer such that $11\leq p\leq n-$ 1. Assume that at any point $x\in M^{n}$ and for any orthonormal basis $\{e_{1},\ldots,e_{n}\}$ of $T_{x}M$ the second fundamental form $\alpha_{f}\colon TM\times TM\to N_{f}M$ satisfies
[…]
Lemma 6.
Let $f\colon M^{4}\to\mathbb{Q}_{c}^{4+m},c\geq 0$, be an isometric immersion of a $4$-dimensional Riemannian manifold $M^{4}$ such that the inequality (italic-$\ast$1) is satisfied at a point $x\in M^{4}$ for $k=2$. Then the following assertions hold at $x$:
[…]
$\pi_{j}\circ A_{\xi}|_{V_{j}}={\langle}\xi,\eta_{j}{\rangle}Id,$
where $Id$ is the identity map on the tangent space at $x$, $V_{1}=\mbox{span}\{e_{1},e_{2}\}$, $V_{2}=\mbox{span}\{e_{3},e_{4}\}$ and $\pi_{j}$ denotes the projection onto $V_{j},j=1,2$.
3 Proofs of the main results
We recall a key result, specifically Proposition 16 from [17], which establishes an estimate for the Bochner-Weitzenböck operator of a submanifold in terms of its second fundamental form. For four-dimensional submanifolds, this proposition can be stated as follows:
Proposition 7.
Let $f\colon M^{4}\to\mathbb{Q}_{c}^{4+m}$ be an isometric immersion of a $4$-dimensional manifold $M^{4}$. The Bochner-Weitzenböck operator $\mathcal{B}^{[2]}$ of $M^{4}$ satisfies the following pointwise inequality
$\mathop{\min_{\omega\in\Omega^{2}(M^{4})}}_{\|\omega\|=1}{\langle}\mathcal{B}^{[2]}\omega,\omega{\rangle}\geq 4c+8H^{2}-S.$ (2)
If equality holds in (2) at a point $x\in M^{4}$, then the shape operator $A_{\xi}(x)$ has at most two distinct eigenvalues, with multiplicities $2$, for every unit vector $\xi\in N_{f}M(x)$.
We now state the following auxiliary results.
Proposition 8.
Let $f\colon M^{4}\to\mathbb{Q}^{4+m}_{c},c\geq 0$, be an isometric immersion of an oriented $4$-dimensional Riemannian manifold. Suppose that $f$ satisfies the inequality (italic-$\ast$1) at any point for $k=2$. Then, the following assertions hold:
$(i)$ The Bochner-Weitzenböck operator $\mathcal{B}^{[2]}$ is nonnegative, and the manifold $M^{4}$ has nonnegative isotropic curvature.
$(ii)$ Suppose that $M^{4}$ is compact and $\beta_{2}(M)\neq 0$. Then, equality holds in (italic-$\ast$1) at every point. Furthermore, at each point $x\in M^{4}$, there exists an oriented orthonormal four-frame $\{e_{1},e_{2},e_{3},e_{4}\}\subset T_{x}M$ such that the second fundamental form $\alpha_{f}$ of $f$ satisfies the following conditions:
$\;\;\;\;\alpha_{11}=\alpha_{22},\alpha_{33}=$ $\alpha_{44},\alpha_{12}=\alpha_{34}=0,\,\|\alpha_{23}\|=\|\alpha_{14}\|,\|\alpha_{24}\|=\|\alpha_{13}\|,$ (3)
${\langle}\alpha_{14}+\alpha_{23}$ $,\alpha_{13}-\alpha_{24}{\rangle}=0,\;{\langle}\alpha_{13}+\alpha_{24},\alpha_{14}-\alpha_{23}{\rangle}=0,$ (4)
$\|\alpha_{13}\|^{2}+\|\alpha_{14}\|^{2}=c+{\langle}\alpha_{11},\alpha_{44}{\rangle}.$ (5)
Here, for simplicity, we denote $\alpha_{ij}=\alpha_{f}(e_{i},e_{j})$.
Proof.
$(i)$ From ($\ast$1), if follows that the right-hand side of inequality (2) in Proposition 7 is nonnegative. Consequently, the Bochner-Weitzenböck operator $\mathcal{B}^{[2]}$ is nonnegative. By Lemma 4, this implies that the manifold $M^{4}$ has nonnegative isotropic curvature.
$(ii)$ By the Hodge theorem, there exists a nontrivial harmonic 2-form $\omega\in\Omega^{2}(M)$. The Bochner-Weitzenböck formula then gives ${\langle}\mathcal{B}^{[2]}\omega,\omega{\rangle}=0$ at any point. Therefore, equality holds in inequality (2), and as a result, Proposition 7 implies that the shape operator, associated with any normal direction has at most two distinct eigenvalues, each with multiplicity $2$.
[…]
$\begin{cases}A_{\alpha}e_{1}=\rho_{\alpha}e_{1}+\kappa_{\alpha}e_{3}+\lambda_{\alpha}e_{4}\\ A_{\alpha}e_{2}=\rho_{\alpha}e_{2}+\mu_{\alpha}e_{3}+\nu_{\alpha}e_{4}\\ A_{\alpha}e_{3}=\kappa_{\alpha}e_{1}+\mu_{\alpha}e_{2}+\sigma_{\alpha}e_{3}\\ A_{\alpha}e_{4}=\lambda_{\alpha}e_{1}+\nu_{\alpha}e_{2}+\sigma_{\alpha}e_{4},\end{cases}$ (6)
where
$\rho_{1}+\sigma_{1}=2H\;\;\text{and}\;\;\rho_{\alpha}+\sigma_{\alpha}=0\;\;\text{for any}\;\;2\leq\alpha\leq m.$
Since each shape operator has at most two distinct eigenvalues, each with multiplicity $2$, it follows from (6) that
$\nu_{\alpha}=\pm\kappa_{\alpha}\;\;\text{and}\;\;\mu_{\alpha}=\mp\lambda_{\alpha}\;\;\text{for any}\;\;1\leq\alpha\leq m.$
From this, (3) follows directly, along with the relations:
${\langle}\alpha_{13},\alpha_{14}{\rangle}+{\langle}\alpha_{23},\alpha_{24}{\rangle}=0,\;{\langle}\alpha_{13},\alpha_{23}{\rangle}+{\langle}\alpha_{14},\alpha_{24}{\rangle}=0.$ (7)
Using (3), (7), and the relation $\alpha_{11}+\alpha_{44}=\mathcal{H}/2$, we deduce that ($\ast$1), which now holds as equality, is equivalent to (5). Furthermore, if (4) is satisfied, the orthonormal four-frame $\{e_{i}\}_{1\leq i\leq 4}$ satisfies all the desired properties.
Suppose now that the frame $\{e_{i}\}_{1\leq i\leq 4}$ does not satisfy (4). Consider instead the orthonormal four-frame $\{\tilde{e}_{i}\}_{1\leq i\leq 4}$, where $\tilde{e}_{i}=R_{\theta}e_{i},\tilde{e}_{j}=R_{\varphi}e_{j}$, for $i=1,2,j=3,4$. Here, $R_{\theta}$ and $R_{\varphi}$ denote rotations on $V_{1}=\mbox{span}\{e_{1},e_{2}\}$ and $V_{2}=\mbox{span}\{e_{3},e_{4}\}$ by angles $\theta$ and $\varphi$, respectively. For simplicity, we set $\tilde{\alpha}_{ij}=\alpha_{f}(\tilde{e}_{i},\tilde{e}_{j})$. Then, using (3), we have
$\tilde{\alpha}_{ii}=\alpha_{ii},\;1\leq i\leq 4,\;\;{\text{and}}\;\;\tilde{\alpha}_{12}=\tilde{\alpha}_{34}=0.$
Since each shape operator has at most two distinct eigenvalues, each with multiplicity $2$, following a similar argument as above, we conclude that the vectors $\tilde{\alpha}_{ij}$, for $1\leq i,j\leq 4$, satisfy (3), (5) and (7). We now claim that the angles $\theta$ and $\varphi$ can be chosen such that (4) is satisfied for the frame $\{\tilde{e}_{i}\}_{1\leq i\leq 4}$. Straightforward computations yield the following relations:
$\tilde{\alpha}_{14}+\tilde{\alpha}_{23}=\cos(\varphi+\theta)(\alpha_{14}+\alpha_{23})+\sin(\varphi+\theta)(\alpha_{24}-\alpha_{13}),$ (8)
$\tilde{\alpha}_{24}-\tilde{\alpha}_{13}=-\sin(\varphi+\theta)(\alpha_{14}+\alpha_{23})+\cos(\varphi+\theta)(\alpha_{24}-\alpha_{13}),$ (9)
$\tilde{\alpha}_{13}+\tilde{\alpha}_{24}=\cos(\varphi-\theta)(\alpha_{13}+\alpha_{24})+\sin(\varphi-\theta)(\alpha_{14}-\alpha_{23}),$ (10)
$\tilde{\alpha}_{14}-\tilde{\alpha}_{23}=-\sin(\varphi-\theta)(\alpha_{13}+\alpha_{24})+\cos(\varphi-\theta)(\alpha_{14}-\alpha_{23}).$ (11)
We can select angles $\sigma_{1}$ and $\sigma_{2}$ such that the following are satisfied
$2\cos\sigma_{1}{\langle}\alpha_{14}+\alpha_{23},\alpha_{24}-\alpha_{13}{\rangle}+\sin\sigma_{2}\big{(}\|\alpha_{13}-\alpha_{24}\|^{2}-\|\alpha_{14}+\alpha_{23}\|^{2}\big{)}=0,$
$0\displaystyle 2\cos\sigma_{2}{\langle}\alpha_{13}+\alpha_{24},\alpha_{14}-\alpha_{23}{\rangle}-\sin\sigma_{2}\big{(}\|\alpha_{14}-\alpha_{23}\|^{2}-\|\alpha_{13}+\alpha_{24}\|^{2}\big{)}=$ 0 .
By choosing $\varphi=(\sigma_{1}+\sigma_{2})/4,\theta=(\sigma_{1}-\sigma_{2})/4$, and applying (8)-(11), it follows that
${\langle}\tilde{\alpha}_{14}+\tilde{\alpha}_{23},\tilde{\alpha}_{13}-\tilde{\alpha}_{14}{\rangle}=0,\;{\langle}\tilde{\alpha}_{13}+\tilde{\alpha}_{24},\tilde{\alpha}_{14}-\tilde{\alpha}_{23}{\rangle}=0.$
This completes the proof of the proposition.∎
Lemma 9.
Let $f\colon M^{4}\to\mathbb{Q}^{4+m}_{c},c\geq 0$, be an isometric immersion of an oriented $4$-dimensional Riemannian manifold $M^{4}$. Suppose that at a point $x\in M^{4}$, there exists an oriented orthonormal four-frame $\{e_{1},e_{2},e_{3},e_{4}\}\subset T_{x}M$ such that the second fundamental form of $f$ at $x$ satisfies the conditions (3)-(5) in part (ii) of Proposition 8. We consider the orthonormal basis $\{\eta_{i}\}_{1\leq i\leq 6}$ of the space of 2-vectors $\Lambda^{2}T_{x}M$, satisfying $\eta_{i}\in\Lambda_{+}^{2}T_{x}M$, $\eta_{i+3}=\ast\eta_{i}$, for $1\leq i\leq 3$, and
$\eta_{1}=\frac{1}{\sqrt{2}}(e_{12}+e_{34}),\eta_{2}$ $=\frac{1}{\sqrt{2}}(e_{13}-e_{24}),\eta_{3}=\frac{1}{\sqrt{2}}(e_{14}+e_{23}),$
where we have set $e_{ij}=e_{i}\wedge e_{j}$. Then the following assertions hold at $x$:
$(i)$ The matrix of the Bochner-Weitzenböck operator $\mathcal{B}^{[2]}=\mathcal{B}_{+}^{[2]}\oplus\mathcal{B}_{-}^{[2]}$ at the point $x$, with respect to the basis $\{\eta_{i}\}_{1\leq i\leq 6}$, is given by the direct sum
$\begin{bmatrix}\mu^{+}_{1}&a^{+}_{1}&a^{+}_{2}\\ a^{+}_{1}&\mu^{+}_{2}&0\\ a^{+}_{2}&0&\mu^{+}_{3}&\!\!\!\!\!\end{bmatrix}\oplus\begin{bmatrix}\mu^{-}_{1}&a^{-}_{1}&a^{-}_{2}\\ a^{-}_{1}&\mu^{-}_{2}&0\\ a^{-}_{2}&0&\mu^{-}_{3}&\!\!\!\!\!\end{bmatrix},$
where
$\mu^{\pm}_{1}=\;$ $\|\alpha_{14}\pm\alpha_{23}\|^{2}+\|\alpha_{13}\mp\alpha_{24}\|^{2},$
$\mu^{\pm}_{2}=\;$ $2c+\|\alpha_{11}\|^{2}+\|\alpha_{44}\|^{2}+\|\alpha_{23}\mp\alpha_{14}\|^{2},$
$\mu^{\pm}_{3}=\;$ $2c+\|\alpha_{11}\|^{2}+\|\alpha_{44}\|^{2}+\|\alpha_{13}\pm\alpha_{24}\|^{2},$
$a^{\pm}_{1}=\;$ ${\langle}\alpha_{23}\pm\alpha_{14},\alpha_{44}-\alpha_{11}{\rangle},\;a^{\pm}_{2}={\langle}\alpha_{24}\mp\alpha_{13},\alpha_{44}-\alpha_{11}{\rangle}.$
$(ii)$ If $\ker\mathcal{B}_{+}^{[2]}\neq 0$ at $x$, then one of the following conditions hold:
$a_{14}+\alpha_{23}=0=\alpha_{13}-\alpha_{24}\;\;{\text{a}nd}\;\;a_{1}^{+}=a_{2}^{+}=0,$ (ii1)
$c=0,\;\alpha_{11}=\alpha_{44}=\alpha_{13}=\alpha_{24}\;\;{\text{a}nd}\;\;\alpha_{14}+\alpha_{23}\neq 0,$ (ii2)
$c=0,\;\alpha_{11}=\alpha_{44}=\alpha_{14}=\alpha_{23}\;\;{\text{a}nd}\;\;\alpha_{13}\neq\alpha_{24}.$ (ii3)
$(iii)$ If $\ker\mathcal{B}_{-}^{[2]}\neq 0$ at $x$, then one of the following conditions hold:
$\alpha_{14}-\alpha_{23}=0=\alpha_{13}+\alpha_{24}\;{\text{a}nd}\;a_{1}^{-}=a_{2}^{-}=0,$ (iii1)
$c=0,\;\alpha_{11}=\alpha_{44}=\alpha_{13}=\alpha_{24}\;\;{\text{a}nd}\;\;\alpha_{14}\neq\alpha_{23},$ (iii2)
$c=0,\;\alpha_{11}=\alpha_{44}=\alpha_{14}=\alpha_{23}\;\;{\text{a}nd}\;\;\alpha_{13}+\alpha_{24}\neq 0.$ (iii3)
Proof.
$(i)$ By a straightforward computation using (2.1), the Gauss equation, and (3)-(5), we obtain the following
$\mathcal{B}^{[2]}(\eta_{1})=$ $\;\big{(}\|\alpha_{14}+\alpha_{23}\|^{2}+\|\alpha_{13}-\alpha_{24}\|^{2}\big{)}\eta_{1}$
$+{\langle}\alpha_{14}+\alpha_{23},\alpha_{44}-\alpha_{11}{\rangle}\eta_{2}+{\langle}\alpha_{24}-\alpha_{13},\alpha_{44}-\alpha_{11}{\rangle}\eta_{3}$
$\mathcal{B}^{[2]}(\eta_{2})=$ $\;{\langle}\alpha_{14}+\alpha_{23},\alpha_{44}-\alpha_{11}{\rangle}\eta_{1}$
$+\big{(}2c+\|\alpha_{11}\|^{2}+\|\alpha_{44}\|^{2}+\|\alpha_{23}-\alpha_{14}\|^{2}\big{)}\eta_{2},$
$\mathcal{B}^{[2]}(\eta_{3})=$ $\;{\langle}\alpha_{24}-\alpha_{13},\alpha_{44}-\alpha_{11}{\rangle}\eta_{1}$
$+\big{(}2c+\|\alpha_{11}\|^{2}+\|\alpha_{44}\|^{2}+\|\alpha_{13}+\alpha_{24}\|^{2}\big{)}\eta_{3},$
and
$\mathcal{B}^{[2]}(\eta_{4})=$ $\;\big{(}\|\alpha_{14}-\alpha_{23}\|^{2}+\|\alpha_{13}+\alpha_{24}\|^{2}\big{)}\eta_{4}$
$+{\langle}\alpha_{23}-\alpha_{14},\alpha_{44}-\alpha_{11}{\rangle}\eta_{5}+{\langle}\alpha_{24}+\alpha_{13},\alpha_{44}-\alpha_{11}{\rangle}\eta_{6}$
$\mathcal{B}^{[2]}(\eta_{5})=$ $\;{\langle}\alpha_{23}-\alpha_{14},\alpha_{44}-\alpha_{11}{\rangle}\eta_{4}$
$+\big{(}2c+\|\alpha_{11}\|^{2}+\|\alpha_{44}\|^{2}+\|\alpha_{23}+\alpha_{14}\|^{2}\big{)}\eta_{5},$
$\mathcal{B}^{[2]}(\eta_{6})=$ $\;{\langle}\alpha_{24}+\alpha_{13},\alpha_{44}-\alpha_{11}{\rangle}\eta_{4}$
$+\big{(}2c+\|\alpha_{11}\|^{2}+\|\alpha_{44}\|^{2}+\|\alpha_{13}-\alpha_{24}\|^{2}\big{)}\eta_{6}.$
This completes the proof of part $(i)$.
$(ii)$ Using part $(i)$, we find that
$\det\mathcal{B}_{+}^{[2]}=\mu^{+}_{1}\mu^{+}_{2}\mu^{+}_{3}-\mu^{+}_{2}(a_{2}^{+})^{2}-\mu^{+}_{3}(a_{1}^{+})^{2}.$
Hence, if $\ker\mathcal{B}_{+}^{[2]}\neq 0$ at $x$, then
$\mu^{+}_{1}\mu^{+}_{2}\mu^{+}_{3}=\mu^{+}_{2}(a_{2}^{+})^{2}+\mu^{+}_{3}(a_{1}^{+})^{2},$
or equivalently
$\mu^{+}_{2}\big{(}\mu^{+}_{3}\|\alpha_{13}-\alpha_{24}\|^{2}-(a_{2}^{+})^{2}\big{)}+\mu^{+}_{3}\big{(}\mu^{+}_{2}\|\alpha_{14}+\alpha_{23}\|^{2}-(a_{1}^{+})^{2}\big{)}=0.$ (12)
On the other hand, by applying the Cauchy-Schwarz inequality and using (5), we obtain
$\mu^{+}_{2}\big{(}\mu^{+}_{3}\|\alpha_{13}-\alpha_{24}\|^{2}$ $-(a_{2}^{+})^{2}\big{)}+\mu^{+}_{3}\big{(}\mu^{+}_{2}\|\alpha_{14}+\alpha_{23}\|^{2}-(a_{1}^{+})^{2}\big{)}$
$\geq$ $\;\mu^{+}_{2}\|\alpha_{13}-\alpha_{24}\|^{2}\big{(}\mu^{+}_{3}-\|\alpha_{44}-\alpha_{11}\|^{2}\big{)}$
$+\mu^{+}_{3}\|\alpha_{14}+\alpha_{23}\|^{2}\big{(}\mu^{+}_{2}-\|\alpha_{44}-\alpha_{11}\|^{2}\big{)}$
$=$ $\;\mu^{+}_{2}\|\alpha_{13}-\alpha_{24}\|^{2}\big{(}2\|\alpha_{13}\|^{2}+2\|\alpha_{14}\|^{2}+\|\alpha_{13}+\alpha_{24}\|^{2}\big{)}$
$+\mu^{+}_{3}\|\alpha_{14}+\alpha_{23}\|^{2}\big{(}2\|\alpha_{13}\|^{2}+2\|\alpha_{14}\|^{2}+\|\alpha_{23}-\alpha_{14}\|^{2}\big{)}.$
The above together with (12), implies that
$\mu^{+}_{2}\|\alpha_{13}-\alpha_{24}\|^{2}\big{(}2\|\alpha_{13}\|^{2}+2\|\alpha_{14}\|^{2}+\|\alpha_{13}+\alpha_{24}\|^{2}\big{)}=0,$
$\mu^{+}_{3}\|\alpha_{14}+\alpha_{23}\|^{2}\big{(}2\|\alpha_{13}\|^{2}+2\|\alpha_{14}\|^{2}+\|\alpha_{23}-\alpha_{14}\|^{2}\big{)}=0.$
This, together with (5), clearly concludes the proof of part $(ii)$. The proof of part $(iii)$ follows similarly and is therefore omitted. ∎
Let $f\colon M^{n}\to\mathbb{Q}_{c}^{n+m}$ be an isometric immersion. We recall that a vector $\eta$ in the normal space $N_{f}M(x)$ is called a Dupin principal normal of the isometric immersion $f$ at a point $x\in M^{n}$ if the tangent subspace
$E_{\eta}(x)=\left\{X\in T_{x}M:\alpha_{f}(X,Y)={\langle}X,Y{\rangle}\eta\;\,\text{for all}\;\,Y\in T_{x}M\right\}$
is at least two dimensional. That dimension is called the multiplicity of $\eta$. The relative nullity subspace $\mathcal{D}_{f}(x)$ of $f$ at a point $x\in M^{n}$ is the kernel of the second fundamental form at this point, namely
$\mathcal{D}_{f}(x)=\left\{X\in T_{x}M:\alpha_{f}(X,Y)=0\;\,\text{for all}\;\,Y\in T_{x}M\right\}.$
We will need the following result.
Proposition 10.
Let $f\colon M^{n}\to\mathbb{Q}_{c}^{n+m},n\geq 4,c\geq 0$, be a substantial isometric immersion of a compact, simply connected, even-dimensional manifold with flat normal bundle. Assume that equality holds in (italic-$\ast$1) for $k=n/2$ at every point, and that there exist two principal normals of multiplicity $k$. Then, $M^{n}$ is isometric to a torus $\mathbb{S}^{k}(r)\times\mathbb{S}^{k}(\sqrt{R^{2}-r^{2}})$ and $f$ is a composition $f=j\circ g$, where $g\colon M^{n}\to\mathbb{S}^{n+1}(R)$ is the standard embedding of the torus $\mathbb{S}^{k}(r)\times\mathbb{S}^{k}(\sqrt{R^{2}-r^{2}})$ into a sphere $\mathbb{S}^{n+1}(R)$, and $j\colon\mathbb{S}^{n+1}(R)\to\mathbb{Q}_{c}^{n+2}$ is an umbilical inclusion, with $R\leq 1$ if $c=1$.
Proof.
Let $\eta_{1},\eta_{2}$ be the principal normals with corresponding distributions $E_{1},E_{2}$. Since $2\mathcal{H}_{f}=\eta_{1}+\eta_{2}$ and $S=k(\|\eta_{1}\|^{2}+\|\eta_{2}\|^{2})$, condition ($\ast$1), which now holds as equality for $k=n/2$, implies
${\langle}\eta_{1},\eta_{2}{\rangle}=-c.$ (13)
Consider the open subset $M^{*}$ of points where $f$ is not umbilical, i.e., where $\eta_{1}\neq\eta_{2}$. On $M^{*}$, the vector fields $\eta_{1}$ and $\eta_{2}$ are smooth Dupin principal normals, with smooth corresponding distributions $E_{1},E_{2}$, each of rank $k$.
The Codazzi equation for $f$ is easily seen to yield
${\langle}\nabla_{X}Y,Z{\rangle}(\eta_{i}-\eta_{j})={\langle}X,Y{\rangle}\nabla_{Z}^{\perp}\eta_{i}\;\,\text{if}\;\,i\neq j,$ (14)
for all $X,Y\in E_{i},Z\in E_{j},1\leq i,j\leq 2$. Using (13) and the fact that $\eta_{1}$ and $\eta_{2}$ are Dupin principal normal vector fields, (14) implies that the following relations hold on $M^{*}$:
${\langle}\nabla_{X}Y,Z{\rangle}(c+\|\eta_{j}\|^{2})=0\;\,\text{for all}\;\,X,Y\in E_{i}\;\,\text{and}\;\,Z\in E_{j},\;i\neq j.$ (15)
Now we distinguish two cases.
Case $c>0$. From (13), it is clear that $M^{*}=M^{n}$. Consequently, it follows from (15) that the distributions $E_{1}$ and $E_{2}$ are totally geodesic on $M^{n}$. Since $M^{n}$ is simply connected, it is well-known that $M^{n}$ is a Riemannian product $M_{1}^{k}\times M_{2}^{k}$ (cf. Theorem $8.2$ in [5]) such that $TM_{i}^{k}=E_{i}$, $i=1,2$. As the second fundamental form of $f$ is adapted to the distributions $E_{1}$ and $E_{2}$, then Theorem $8.4$ and Corollary $8.6$ in [5] imply that the submanifold is an extrinsic product of isometric immersions, each of which is totally umbilical. Therefore, the submanifold is a torus $\mathbb{S}^{k}(r)\times\mathbb{S}^{k}(\sqrt{R^{2}-r^{2}})$ in a sphere $\mathbb{S}^{n+1}(R)\subset\mathbb{S}^{n+2}$.
Case $c=0$. Consider the open subset
$M_{+}=\{x\in M^{*}:\eta_{1}(x)\neq 0\neq\eta_{2}(x)\}.$
Suppose that $M_{+}$ is nonempty. From (15), it follows that that the distributions $E_{1}$ and $E_{2}$ are totally geodesic on $M_{+}$. It is well-known that $M_{+}$ is locally a Riemannian product $M_{1}^{k}\times M_{2}^{k}$ (cf. Theorem $8.2$ in [5]) such that $TM_{i}^{k}=E_{i}$, $i=1,2$. Since the second fundamental form of $f$ is adapted to the distributions $E_{1}$ and $E_{2}$, Theorem $8.4$ and Corollary $8.6$ in [5] imply that the submanifold is locally an extrinsic product of isometric immersions, each of which is totally umbilical. Consequently, the submanifold is locally, on $M_{+}$, a torus $\mathbb{S}^{k}(r)\times\mathbb{S}^{k}(\sqrt{R^{2}-r^{2}})$ in a sphere $\mathbb{S}^{n+1}(R)\subset\mathbb{R}^{n+2}$.
Now, suppose that the interior ${\rm int}(M_{0})$ of the subset $M_{0}=M^{*}\smallsetminus M_{+}$ is nonempty. Assume that $\eta_{2}=0$ on a connected component $U$ of ${\rm int}(M_{0})$. Then $f$ has constant index of relative nullity $k$ on $U$. From (15), it follows that the relative nullity distribution $\mathcal{D}_{f}=E_{2}$ is totally geodesic. Let $C_{T}\colon\mathcal{D}_{f}^{\perp}\to\mathcal{D}_{f}^{\perp}$ be the associated splitting tensor for any $T\in\mathcal{D}_{f}$ (see [5, p. 186]). Since
$\alpha_{f}(X,Y)={\langle}X,Y{\rangle}\eta_{1}\;\;{\text{for all}}\;\;X\in\mathcal{X}(M^{n}),Y\in\mathcal{D}_{f}^{\perp},$
the Codazzi equation
$(\nabla_{X}^{\perp}\alpha_{f})(Y,T)=(\nabla_{T}^{\perp}\alpha_{f})(X,Y)\;\;{\text{for all}}\;\;X,Y\in\mathcal{D}_{f}^{\perp},T\in\mathcal{D}_{f}$
implies that $C_{T}={\langle}\nabla\log\|\eta_{1}\|,T{\rangle}Id$, where $Id$ is the identity map on the conullity distribution $\mathcal{D}_{f}^{\perp}$. This shows that the conullity distribution is umbilical, and consequently, integrable.
Let $\varSigma^{k}$ be a leaf of $\mathcal{D}_{f}^{\perp}$. By Proposition 7.6 in [5], $f$ is locally a generalized cone over an isometric immersion $g\colon\varSigma^{k}\to\mathbb{Q}_{\tilde{c}}^{k+m}$ such that $f\circ j_{\varSigma}=i\circ g$, where $j_{\varSigma}\colon\varSigma^{k}\to M^{n}$ is the inclusion and $i\colon\mathbb{Q}_{\tilde{c}}^{k+m}\to\mathbb{R}^{n+m}$ is an umbilical inclusion. The second fundamental form of the immersion $f\circ j$ is given by
$\alpha_{f\circ j}(X,Y)={\langle}X,T{\rangle}\big{(}\eta_{1}+f_{*}\nabla\log\|\eta_{1}\|\big{)}\;\;{\text{for all}}\;\;X,Y\in\mathcal{X}(\varSigma^{k}).$
Thus, $f\circ j_{\varSigma}$ is umbilical and $g(\varSigma^{k})$ is a sphere $\mathbb{S}^{k}(r)$ centered at a point $x_{0}\in\mathbb{R}^{n+m}$ with radius $r=1/\sqrt{\lambda^{2}+\|\nabla\log\|\eta_{1}\|\|^{2}}$. If the umbilical submanifold $\mathbb{Q}_{\tilde{c}}^{k+m}$ is totally geodesic in $\mathbb{R}^{n+m}$, then the submanifold $f$ is a $k$-cylinder over the sphere $\mathbb{S}^{k}(r)$. Alternatively, if $\mathbb{Q}_{\tilde{c}}^{k+m}$ is a sphere centered at a point $\tilde{x}_{0}\in\mathbb{R}^{n+m}$, then $f$ is a $(k-1)$-cylinder over a submanifold $N^{k+1}$, which is a cone over the sphere $\mathbb{S}^{k}(r)$ with its vertex at $\tilde{x}_{0}\neq x_{0}$.
Consequently, on ${\rm int}(M_{0})$ the submanifold is locally a $k$-cylinder in $\mathbb{R}^{n+1}$ over a sphere $\mathbb{S}^{k}(r)$, or a $(k-1)$-cylinder over a submanifold $N^{k+1}$ in $\mathbb{R}^{k+2}$, which is a cone over a sphere $\mathbb{S}^{k}(r)$.
In the later case, $f$ is given locally on $\mathrm{int}(M_{0})$ by
$f(x,w)=j(x)+w,\;\;(x,w)\in N_{i}\mathbb{S}^{k+m}(R),$
where the normal bundle $N_{i}\mathbb{S}^{k+m}(R)$ of the inclusion $i\colon\mathbb{S}^{k+m}(R)\to\mathbb{R}^{n+m}$ is regarded as subbundle of $N_{i\circ j}\mathbb{S}^{k}(r)$. Here, $j\colon\mathbb{S}^{k}(r)\to\mathbb{S}^{k+m}(R)$ is an umbilical inclusion with $r<R$. Equivalently, $f$ is locally parametrized by
$f(x,t_{0},t_{1},\dots,t_{k-1})=t_{0}j(x)+\sum_{i=1}^{k-1}t_{\ell}v_{\ell},\;t_{0}>0,\;x\in\mathbb{S}^{k}(r),$
where $\mathbb{S}^{k+m}(R)\subset\mathbb{R}^{k+m+1}$, $\mathbb{R}^{n+m}=\mathbb{R}^{k+m+1}\oplus\mathbb{R}^{k-1}$, and $\{v_{\ell}\}_{1\leq\ell\leq k-1}$ is an orthonormal basis of $\mathbb{R}^{k-1}$. The Laplacian operator $\Delta_{M}$ of $M^{n}$ is given by
$\Delta_{M}=\frac{k}{t_{0}}\frac{\partial}{\partial t_{0}}+\frac{\partial^{2}}{\partial t_{0}^{2}}+\frac{1}{t^{2}_{0}}\Delta_{\mathbb{S}^{k}(r)}+\sum_{\ell=1}^{k-1}\frac{\partial^{2}}{\partial t^{2}_{\ell}}.$
Since $S=a/t^{2}_{0}$, where $a$ is a positive constant, it follows that
$\Delta_{M}S=-\frac{a(n-6)}{t_{0}^{4}}.$
Consequently, either $\Delta_{M}S\geq 0$ or $\Delta_{M}S\leq 0$ on ${\rm int}(M_{0})$, depending on the dimension $n$.
Since $S$ is locally constant on $M_{+}$, continuity implies that $S$ is either superharmonic or subharmonic function on $M^{n}$. By the maximum principle, it follows that $S$ is a positive constant on $M^{n}$. Consequently, $M^{*}=M^{n}$, and therefore $M^{n}=M_{+}\cup M_{0}$.
Thus, on $M_{0}$, one of the Dupin principal normal vector fields vanishes, while the other has constant length $S/k$. On the other hand, both principal normals have constant and positive length on each connected component of $M_{+}$. By continuity and the compactness of $M^{n}$, it follows that $M_{+}=M^{n}$. Consequently, using a similar argument as in Case $c>0$, we conclude that $M^{n}$ is isometric to a torus $\mathbb{S}^{k}(r)\times\mathbb{S}^{k}(\sqrt{R^{2}-r^{2}})$, and $f$ is the standard embedding into a sphere $\mathbb{S}^{n+1}(R)\subset\mathbb{R}^{n+2}$. ∎
Proof of Theorem 2: Part $(i)$ of Proposition 8 implies that $M^{4}$ has nonnegative isotropic curvature and that the Bochner-Weitzenböck operator is nonnegative at any point. The proof of the theorem is divided into three cases.
Case I. We begin by proving the theorem for simply connected submanifolds. Since $M^{4}$ has nonnegative isotropic curvature, Theorem 4.10 in [15] implies that one of the following holds:
  1. (a)
    $M^{4}$ carries a metric of positive isotropic curvature.
  2. (b)
    $M^{4}$ is diffeomorphic to a product $\mathbb{S}^{2}\times\varSigma^{2}$, where $\varSigma^{2}$ is a compact surface.
  3. (c)
    $M^{4}$ is a Kähler manifold biholomorphic to $\mathord{\mathbb{C}}\mathord{\mathbb{P}}^{2}$.
We analyse each case separately as follows.
Case (a). We assert that the manifold $M^{4}$ is diffeomorphic to $\mathbb{S}^{4}$. Given that $M^{4}$ is simply connected by assumption, the main result in [14] implies that $M^{4}$ is homeomorphic to $\mathbb{S}^{4}$. Furthermore, $M^{4}$ is locally irreducible. If this were not the case, then Theorem 3.1 in [15] would imply that $M^{4}$ is isometric to a Riemannian product of two compact surfaces, which leads to a contradiction. Hence, $M^{4}$ is locally irreducible, and by Theorem 2 in [3], one of the following cases must hold:
  1. (i)
    $M^{4}$ is diffeomorphic to a spherical space form.
  2. (ii)
    $M^{4}$ is a Kähler manifold biholomorphic to $\mathord{\mathbb{C}}\mathord{\mathbb{P}}^{2}$.
  3. (iii)
    $\;M^{4}$ is isometric to a compact symmetric space.
Since $M^{4}$ is homeomorphic to $\mathbb{S}^{4}$, case (ii) above is ruled out. Clearly, in case (i), the manifold must be diffeomorphic to $\mathbb{S}^{4}$. Furthermore, the only 4-dimensional compact symmetric spaces are the round spheres, the product of two 2-dimensional spheres, or the complex or the quaternionic projective space $\mathbb{H}P^{1}$. Consequently, in case (iii), the manifold $M^{4}$ must be diffeomorphic to $\mathbb{S}^{4}$.
Case (b). Since $M^{4}$ is simply connected, the surface $\varSigma^{2}$ must be diffeomorphic to $\mathbb{S}^{2}$, implying that $M^{4}$ is diffeomorphic to $\mathbb{S}^{2}\times\mathbb{S}^{2}$. Consequently, $\beta_{\pm}(M)=1$. It then follows from part $(ii)$ of Proposition 8 that equality holds in ($\ast$1). Moreover, at each point $x\in M^{4}$, there exists an oriented orthonormal four-frame $\{e_{1},e_{2},e_{3},e_{4}\}\subset T_{x}M$ such that the second fundamental form $\alpha_{f}$ of $f$ satisfies the conditions (3)-(5).
In addition, there exist a nontrivial self-dual harmonic 2-form $\omega_{+}$ and a nontrivial anti-self-dual harmonic 2-form $\omega_{-}$. By the Bochner-Weitzenböck formula, both $\omega_{+}$ and $\omega_{-}$ are parallel and ${\langle}\mathcal{B}^{[2]}(\omega_{\pm}),\omega_{\pm}{\rangle}=0$ at any point. Consequently, $\ker\mathcal{B}_{+}^{[2]}\neq 0$ and $\ker\mathcal{B}_{-}^{[2]}\neq 0$ at any point. Using Lemma (9) and (3)-(5), we deduce that only cases $(ii1)$ and $(iii1)$ in that lemma can occur at any point. Furthermore, applying (3)-(5) once more, it follows that four-frame $\{e_{1},e_{2},e_{3},e_{4}\}\subset T_{x}M$ diagonalizes the second fundamental form at any $x\in M^{4}$. As a result, $f$ has flat normal bundle. Moreover, $\alpha_{11}$ and $\alpha_{44}$ are principal normals of $f$ with multiplicity $2$ satisfying ${\langle}\alpha_{11},\alpha_{44}{\rangle}=-c$. It then it follows from Proposition 10 that the submanifold is as described in part $(iiia)$ of the theorem.
Case (c). In this case, either $\beta_{+}(M)=1$ and $\beta_{-}(M)=0$, or $\beta_{+}(M)=0$ and $\beta_{-}(M)=1$. We will only treat the former case, as the latter one can be handled in a similar manner. It follows then from part $(ii)$ of Proposition 8 that equality holds in ($\ast$1). Furthermore, at each point $x\in M^{4}$ there exists an oriented orthonormal four-frame $\{e_{1},e_{2},e_{3},e_{4}\}\subset T_{x}M$ such that the second fundamental form $\alpha_{f}$ of $f$ satisfies the conditions (3)-(5).
Clearly, there exists a nontrivial self-dual harmonic 2-form $\omega_{+}$. By the Bochner-Weitzenböck formula, the form $\omega_{+}$ is parallel, and ${\langle}\mathcal{B}^{[2]}(\omega_{+}),\omega_{+}{\rangle}=0$ at any point. Hence, $\ker\mathcal{B}_{+}^{[2]}\neq 0$ at any point, and part $(ii)$ of Lemma (9) applies. Using (3) and (5), we conclude that only case $(ii1)$ in that lemma can occur at any point. Thus, we have
$\alpha_{14}+\alpha_{23}=\alpha_{13}-\alpha_{24}=0.$ (16)
We assert that $\mu^{+}_{2}\mu^{+}_{3}>0$ at some point. Since $c\geq 0$, if $\mu^{+}_{2}\mu^{+}_{3}=0$ at a point $x$, it follows that $c=0$. Using (3)-(5), we deduce that $f$ must be totally geodesic at $x$. Therefore, there must exist a point $x_{0}$ where $\mu^{+}_{2}\mu^{+}_{3}>0$. From part $(i)$ of Lemma 9, it follows that the kernel of $\mathcal{B}_{+}^{[2]}$ at $x_{0}$ is spanned by the vector $e_{1}\wedge e_{2}+e_{3}\wedge e_{4}$.
Let $X$ be the dual to the self-dual form $\omega_{+}$. Since this form is parallel, we may assume, after possibly multiplying by a constant such that $\|\omega_{+}\|=\sqrt{2}$, that at the point $x_{0}$ we have $X_{x_{0}}=e_{1}\wedge e_{2}+e_{3}\wedge e_{4}.$ Now, consider the almost complex structure $J_{x_{0}}\colon T_{x_{0}}M\to T_{x_{0}}M$ given by $J_{x_{0}}e_{1}=e_{2}$ and $J_{x_{0}}e_{3}=e_{4}$. Then, we have
$\omega_{+}(v,w)={\langle}v,J_{x_{0}}w{\rangle}\;\;\text{for any}\>\>v,w\in T_{x_{0}}M.$
Moreover, we define the skew-symmetric endomorphism $J$ of the tangent bundle of $M^{4}$ such that
$\omega_{+}(X,Y)={\langle}X,JY{\rangle}\;\;\text{for any}\>\>X,Y\in\mathcal{X}(M).$
Clearly, $J$ is parallel because $\omega_{+}$ is parallel. Now we claim that $J$ is orthogonal, i.e., $\|J_{x}v\|=\|v\|$ for any point $x\in M^{4}$ and any $v\in T_{x}M$. Indeed, let $V$ be a parallel vector field along a curve $c\colon[0,1]\to M$ such that $c(0)=x,c(1)=x_{0}$ and $V(0)=v$. Obviously, the vector field $W=JV$ is also parallel along $c$. Using that $J_{x_{0}}$ is orthogonal we have
$\|J_{x}v\|=\|W(0)\|=\|W(1)\|=\|J_{x_{0}}V(1)\|=\|V(1)\|=\|V(0)\|=\|v\|,$
which proves the claim. Since $J$ is both skew-symmetric and orthogonal, we have that $J$ is an almost complex structure that is also parallel. Hence, the triple $(M^{4},{\langle}\cdot,\cdot{\rangle},J)$ is a Kähler manifold.
In case $c>0$, we have that $\mu_{1}^{+}=0$ and $\mu_{2}^{+}\mu_{3}^{+}>0$ at any point. Hence, $X_{x}=e_{1}\wedge e_{2}+e_{3}\wedge e_{4}$ at any point $x$. Thus, we have that $Je_{1}=e_{2}$ and $Je_{3}=e_{4}$ at any point. From (16), it follows that the second fundamental form of the submanifold satisfies
$\alpha_{f}(JX,JY)=\alpha_{f}(X,Y)\;\;\text{for all}\;\;X,Y\in\mathcal{X}(M).$ (17)
In case $c=0$, the above argument applies to the open subset of points where $f$ is not totally geodesic. By continuity, (17) holds everywhere.
Immersions satisfying condition (17) have parallel second fundamental form (see [7, Th. 4]). Moreover, under the immersion $f$, each geodesic of $M$ is mapped into a plane circle. Such submanifolds were classified in [7] and [18]. From this classification, it follows that the submanifold is as described in part $(iiib)$ of the theorem, thereby completing the proof for the case where the manifold is simply connected.
Case II. Now suppose that the fundamental group of $M^{4}$ is finite. Then we claim that $M^{4}$ is simply connected. Consider the universal covering $\pi\colon\tilde{M}^{4}\to M^{4}$. Since the fundamental group of $M^{4}$ is finite, $\tilde{M}^{4}$ must be compact. Moreover, the isometric immersion $\tilde{f}=f\circ\pi$ satisfies ($\ast$1). Therefore, from Case I, we conclude that either $\tilde{M}^{4}$ is diffeomorphic to $\mathbb{S}^{4}$, or the submanifold $\tilde{f}$ is as described in parts $(i)$ or $(ii)$ of the theorem.
Assume first that $\tilde{M}^{4}$ is diffeomorphic to $\mathbb{S}^{4}$. Arguing as in the proof of Case I, we conclude that $M^{4}$ is locally irreducible. Then, by [3, Th. 2], we have that $M^{4}$ is either diffeomorphic to a spherical space form or isometric to a compact symmetric space. In the former case, it is clear that the manifold $M^{4}$ must be diffeomorphic to $\mathbb{S}^{4}$. In the latter case, since the only 4-dimensional compact symmetric spaces are the round spheres, the product of two 2-dimensional spheres, and the complex or the quaternionic projective space $\mathbb{H}P^{1}$, it follows that $M^{4}$ must also be diffeomorphic to $\mathbb{S}^{4}$.
Now, suppose that the submanifold $\tilde{f}$ is as described in parts $(i)$ or $(ii)$ of the theorem. In either case, the covering map $\pi\colon\tilde{M}^{4}\to M^{4}$ must be a diffeomorphism. This observation completes the proof of the theorem in this case.
Case III. Finally, suppose that the fundamental group $\pi_{1}(M)$ of $M^{4}$ is infinite. We first claim that $M^{4}$ is locally reducible. If this were not the case, then by Theorem 2 in [3], it would follow that the universal cover $\tilde{M}^{4}$ is diffeomorphic to a spherical space form, biholomorphic to $\mathord{\mathbb{C}}\mathord{\mathbb{P}}^{2}$, or isometric to a compact symmetric space. Each of these possibilities contradicts the fact that $\tilde{M}^{4}$ is not compact. Therefore, $M^{4}$ must be locally reducible. From this, it follows by Theorem 3.1 in [15] that the universal cover $\tilde{M}^{4}$ is isometric to one of the following:
  1. (i)
    $(\mathbb{R}^{k},g_{0})\times(N_{1}^{n_{1}},g_{1})\times(N_{2}^{n_{2}},g_{2})$, where $k\geq 0,g_{0}$ is the flat Euclidean metric, and either $n_{i}=2$ and $N_{i}=\mathbb{S}^{2}$ has nonnegative Gaussian curvature, or else $n_{i}=3$ and $N_{i}$ is compact with nonnegative Ricci curvature.
  2. (ii)
    $(\varSigma^{2},g_{\varSigma})\times(N^{2},g_{N})$, where $\varSigma^{2}$ is a surface whose Gaussian curvature $K_{\varSigma}$ is negative at some point, and $N^{2}$ is a compact surface with positive Gaussian curvature $K_{N}$.
First, we claim that case (ii) above cannot occur. Suppose, to the contrary, that it does. Clearly the immersion $\tilde{f}=f\circ\pi$ satisfies ($\ast$1), where $\pi\colon\tilde{M}^{4}\to M^{4}$ is the covering map. We consider the isometric immersions $f_{\varSigma}=\tilde{f}\circ i_{\varSigma}$ and $f_{N}=\tilde{f}\circ i_{N}$, where $i_{\varSigma}\colon\varSigma\to\tilde{M}$ and $i_{N}\colon N\to\tilde{M}$ are the inclusion maps. Since both immersions $i_{\varSigma},i_{N}$ are totally geodesic, it follows that the second fundamental form of $f_{\varSigma}$ and $f_{N}$ are given, respectively, by
$\alpha_{f_{\varSigma}}(X,Y)$ $=\alpha_{\tilde{f}}(i_{\varSigma_{*}}X,i_{\varSigma_{*}}Y),\;X,Y\in T\varSigma,$
$\alpha_{f_{N}}(V,W)$ $=\alpha_{\tilde{f}}(i_{N_{*}}V,i_{N_{*}}W),\;V,W\in TN.$
Given that $\tilde{M}^{4}$ is isometric to the Riemannian product $(\varSigma^{2},g_{\varSigma})\times(N^{2},g_{N})$, it follows from the above that
$S_{\tilde{f}}$ $=S_{f_{\varSigma}}+S_{f_{N}}+\sum_{i=1,2}\sum_{j=3,4}\|\alpha_{\tilde{f}}(i_{\varSigma_{*}}e_{i},i_{N_{*}}e_{j})\|^{2},$ (18)
$4\tilde{H}^{2}$ $=H^{2}_{f_{\varSigma}}+H^{2}_{f_{N}}+2{\langle}\mathcal{H}_{f_{\varSigma}},\mathcal{H}_{f_{N}}{\rangle},$ (19)
where $\{e_{i}\}_{1\leq i\leq 4}$ is an orthonormal frame such that $e_{1},e_{2}\in T\varSigma$, $e_{3},e_{4}\in TN$, and $\tilde{H}$ is the mean curvature of the immersion $\tilde{f}$. The squared lengths of the traceless parts $\Phi_{f_{\varSigma}}$ and $\Phi_{f_{N}}$ of the second fundamental forms of the immersions $f_{\varSigma}$ and $f_{N}$ are given by
$\|\Phi_{f_{\varSigma}}\|^{2}=S_{f_{\varSigma}}-4H^{2}_{f_{\varSigma}}\;{\text{and}}\;\|\Phi_{f_{N}}\|^{2}=S_{f_{N}}-4H^{2}_{f_{N}},$
respectively. Taking into account (18), (19) and the above, the inequality ($\ast$1) for $\tilde{f}$ can be equivalently written as
$\|\Phi_{f_{\varSigma}}\|^{2}+\|\Phi_{f_{N}}\|^{2}+2\sum_{i=1,2}\sum_{j=3,4}\|\alpha_{\tilde{f}}(i_{\varSigma_{*}}e_{i},i_{N_{*}}e_{j})\|^{2}\leq 4\big{(}c+{\langle}\mathcal{H}_{f_{\varSigma}},\mathcal{H}_{f_{N}}{\rangle}\big{)}.$ (20)
On the other hand, from the Gauss equation and ($\ast$1), it follows that the scalar curvature $\tau$ of $\tilde{M}^{4}$ satisfies $\tau\geq 4c+8\tilde{H}^{2}$. Since $\tau=2(K_{\varSigma}+K_{N})$, we then obtain the inequality
$K_{\varSigma}+K_{N}\geq 4(c+\tilde{H}^{2}).$ (21)
Using (19) and the Gauss equation for both $f_{\varSigma}$ and $f_{N}$, we find that (21) is written as
$\|\Phi_{f_{\varSigma}}\|^{2}+\|\Phi_{f_{N}}\|^{2}+4\big{(}c+{\langle}\mathcal{H}_{f_{\varSigma}},\mathcal{H}_{f_{N}}{\rangle}\big{)}\leq 0.$ (22)
Then, from (20) and (22) we obtain $\Phi_{f_{\varSigma}}=0$. Hence, the immersion $f_{\varSigma}$ it totally umbilical, and consequently $K_{\varSigma}\geq 0$, which is clearly a contradiction.
Thus, only case (i) holds. Clearly $k\neq 3$. We now claim that $k=1$. If $k=4$, then $M^{4}$ is flat, and by the Gauss equation, we have $S=12c+16H^{2}$. This, together with ($\ast$1), implies that $c=0$ and $f$ is minimal, leading to a contradiction. If $k=2$, then the curvature operator of $M^{4}$ is nonnegative. Then it follows from Theorem 1.3 in [8] that $M^{4}$ is diffeomorphic to a quotient of one of the spaces $\mathbb{S}^{4}$, $\mathord{\mathbb{C}}\mathord{\mathbb{P}}^{4}$, $\mathbb{R}^{1}\times\mathbb{S}^{3}$, $\mathbb{S}^{2}\times\mathbb{S}^{2}$, $\mathbb{R}^{2}\times\mathbb{S}^{2}$, or $\mathbb{R}^{4}$ by a finite group of fixed-point-free isometries in the standard metric. This contradicts the fact that $\pi_{1}(M)$ is infinite. Therefore, $k=1$ and $\tilde{M}$ splits isometrically as ×$\mathbb{R}\times N^{3}$, where $N^{3}$ is compact, simply connected, and it has nonnegative Ricci curvature. From Theorem 1.2 in [8], we then conclude that $N^{3}$ is diffeomorphic to $\mathbb{S}^{3}$.∎
Proof of Corollary 3: Let $n\geq 5$, and suppose that the submanifold $f$ is not as described in parts $(i)$-$(iii)$ of Theorem 1. Then, the homology groups of $M^{n}$ satisfy
$H_{p}(M^{n};\mathbb{Z})=0\;\,\text{for}\;\,2\leq p\leq n-2,\;\,\text{and}\;\,H_{1}(M^{n};\mathbb{Z})\cong\mathbb{Z}^{\beta_{1}(M)}.$
Since $M^{n}$ is oriented, we have $H^{n}(M^{n};\mathbb{Z})\cong\mathbb{Z}$. Applying the universal coefficient theorem, we deduce that $H_{n-1}(M^{n};\mathbb{Z})$ is torsion-free. By Poincaré duality, it follows that $H_{n-1}(M^{n};\mathbb{Z})\cong\mathbb{Z}^{\beta_{1}(M)}$.
For the case $n=4$, the result follows directly from Theorem 2.∎
3.1 Examples of submanifolds satisfying condition ($\ast$1)
We now present a method for constructing geometrically distinct submanifolds that are diffeomorphic to either the sphere $\mathbb{S}^{n}$ or the torus $\mathbb{S}^{1}\times\mathbb{S}^{n-1}$, while also satisfying ($\ast$1) for $c=0$.
Proposition 11.
Let $g\colon N^{n-1}\to\mathbb{R}^{m_{1}},n\geq 4$, be an isometric immersion of a manifold $N^{n-1}$ satisfying condition (italic-$\ast$1) for $k=\ell-1$, with $22\leq\ell\leq n-$ 2, as strict inequality at any point, namely,
$S_{g}<a(n-1,\ell-1,H_{g},0).$
Consider any closed unit-speed curve $\gamma\colon\mathbb{S}^{1}\to\mathbb{R}^{m_{2}}$, whose first curvature $\kappa_{1}$ satisfies the inequality
$\kappa_{1}^{2}\leq\frac{n-\ell}{n-\ell-1}\Big{(}\min\big{(}a(n-1,\ell-1,H_{g},0)-S_{g}\big{)}\Big{)}.$ (23)
Then, the product immersion $f=\gamma\times g\colon\mathbb{S}^{1}\times N^{n-1}\to\mathbb{R}^{m_{1}+m_{2}}$ satisfies condition (italic-$\ast$1) for $k=\ell$.
Proof.
The squared length $S_{f}$ of the second fundamental form, and the mean curvature $H_{f}$ of the product immersion $f$ are given by
$S_{f}=\kappa_{1}^{2}+S_{g}\;\;{\text{and}}\;\;n^{2}H^{2}_{f}=\kappa_{1}^{2}+(n-1)^{2}H^{2}_{g}.$
From these expressions, it follows that condition ($\ast$1) for $f$, with $k=\ell$ and $c=0$, is equivalent to the inequality (23), thereby completing the proof.∎
Next, we prove that there exist many isometric immersions $g\colon N^{n-1}\to\mathbb{R}^{m}$ that satisfy the conditions required in Proposition 11. Specifically, we provide geometrically distinct immersions of $\mathbb{S}^{n}$ into the Euclidean space $\mathbb{R}^{n+1}$ that satisfy ($\ast$1) for any $n\geq 3$ and $11\leq k\leq n-$ 1, either as a strict inequality or otherwise.
Proposition 12.
Let $f\colon M^{n}\to\mathbb{R}^{n+1},n\geq 3$, be an ovaloid in the Euclidean space with principal curvatures $0<\lambda_{1}\leq\dots\leq\lambda_{n}$. If
$\max\lambda_{n}\leq\min\lambda_{1}\Big{(}{\frac{n}{n-k}}\Big{)}^{1/2}$ (24)
for some $11\leq k\leq n-$ 1, then $f$ satisfies condition (italic-$\ast$1). If the inequality (24) is strict, then (italic-$\ast$1) is also strict.
Proof.
Since $a(n,k,H,0)=n^{2}H^{2}/(n-k)$, the proof follows directly from the inequalities $S\leq n\lambda^{2}_{n}$ and $H\geq\lambda_{1}$.∎
A large class of ellipsoids satisfies (24). Consider, for instance, the ellipsoid in $\mathbb{R}^{n+1}$ defined by
$\frac{x_{1}^{2}}{a_{1}^{2}}+\cdots+\frac{x_{n+1}^{2}}{a_{n+1}^{2}}=1,$
where $0<a_{1}\leq\dots\leq a_{n+1}$. A straightforward computation shows that the minimum and the maximum of the principal curvatures of the ellipsoid are $a_{1}/a_{n+1}^{2}$ and $a_{n+1}/a_{1}^{2}$, respectively. It follows that condition (24) is satisfied if
$a_{n+1}\leq a_{1}\big{(}{\frac{n}{n-k}}\big{)}^{1/6}.$
By the classical Hadamard theorem, any ovaloid in $\mathbb{R}^{n+1}$ is diffeomorphic to $\mathbb{S}^{n}$. Consequently, from Propositions 11 and 12, we derive geometrically distinct isometric immersions of manifolds that are diffeomorphic to either the sphere $\mathbb{S}^{n}$ or the torus $\mathbb{S}^{1}\times\mathbb{S}^{n-1},n\geq 4$, which also satisfy ($\ast$1) for $c=0$ and $k=2$. This establishes the optimality of our results.
Example 13.
Consider a rotational hypersurface in $\mathbb{R}^{n+1},n\geq 4$, obtained by rotating a curve $c$ in the $x_{1}x_{2}$-plane around the $x_{1}$-axis. Suppose the curve is given as a graph of a positive function $u=u(x_{1})$. By adjusting the orientation, the rotational hypersurface has two principal curvatures given by
$\lambda=\frac{1}{u(1+(u^{\prime})^{2})^{1/2}},\;\mu=-\frac{u^{\prime\prime}}{(1+(u^{\prime})^{2})^{3/2}},$
with $\lambda$ having multiplicity $n-1$. The hypersurface satisfies the condition ($\ast$1) for $k=2$ as strict inequality if
$(n-1)\lambda^{2}+2(n-1)\lambda\mu-(n-3)\mu^{2}>0,$
or equivalently,
$-a_{n}(1+(u^{\prime})^{2})<uu^{\prime\prime}<b_{n}(1+(u^{\prime})^{2}),$
where
$a_{n}=\frac{\sqrt{2(n-1)(n-2)}+n-1}{n-3},\;b_{n}=\frac{\sqrt{2(n-1)(n-2)}-(n-1)}{n-3}.$
By connecting such hypersurfaces to spheres as in [13, Example 3.3], we obtain compact hypersurfaces that satisfy the condition ($\ast$1) for $k=2$, with the first Betti number being any nonnegative integer.