### v2

# Geometric and topological rigidity of pinched submanifolds II

Abstract
We continue the study of the geometry and topology of compact submanifolds of arbitrary codimension in space forms that satisfy a pinching condition involving the length of the second fundamental form and the mean curvature. Our primary focus is on four-dimensional submanifolds, where both the results obtained and the methods employed differ substantially and are considerably more intricate than in higher dimensions. This study relies critically on concepts from four-dimensional geometry, the theory of Riemannian manifolds with nonnegative isotropic curvature, and the Bochner technique, each playing an essential role. The results are sharp and extend previous work by several authors, without imposing additional assumptions on either the mean curvature or the fundamental group of the submanifold.
footnotetext: 2020 Mathematics Subject Classification. 53C40, 53C42.footnotetext: Keywords and phrases. Homology groups, pinching, length of the second fundamental form, sectional, isotropic and mean curvature, Bochner-Weitzenböck operator.
1 Introduction
A fundamental problem in differential geometry is to understand the interplay between the geometry and topology of Riemannian manifolds. In the context of submanifold theory, it is particularly interesting to investigate how the geometry and topology of submanifolds in space forms are influenced by pinching conditions involving either intrinsic or extrinsic curvature invariants.
For minimal submanifolds of spheres with a sufficiently pinched second fundamental form, a fundamental result was first established by Simons in his seminal work [26]. Later, Chern, do Carmo, and Kobayashi [4] proved a celebrated rigidity theorem. Their contributions have since inspired a wealth of important developments in the study of pinching conditions, including, for example, [1, 11, 14, 20, 25, 27, 32].
In our previous paper [28], we studied the geometric and topological rigidity of submanifolds $f\colon M^{n}\to\mathbb{Q}_{c}^{n+m}$ satisfying the pointwise pinching condition
$S\leq a(n,k,H,c)$ ($\ast$)
where $S$ denotes the squared length of the second fundamental form $\alpha_{f}\colon TM\times TM\to N_{f}M$, taking values in the normal bundle $N_{f}M$. The mean curvature is defined as $H=\|\mathcal{H}\|$, the norm of the mean curvature vector field
$\mathcal{H}=\frac{1}{n}\,\mathrm{tr}\,\alpha_{f},$
where $\mathrm{tr}$ denotes the trace. Here $k$ is an integer with $1\leq k\leq n/2$, and the function $a$ is defined as
$a(n,k,t,c)=nc+\frac{n^{3}t^{2}}{2k(n-k)}-\frac{n|n-2k|}{2k(n-k)}t\sqrt{n^{2}t^{2}+4ck(n-k)},\;t,c\geq 0.$
The ambient space $\mathbb{Q}_{c}^{n+m}$ is the $(n+m)$-dimensional, complete, simply connected space form of constant curvature $c$. For simplicity, we restrict to the cases $c\in\{0,1\}$, unless stated otherwise. Accordingly, $\mathbb{Q}_{c}^{n+m}$ is either the Euclidean space $\mathbb{R}^{n+m}$ when $c=0$, or the unit sphere $\mathbb{S}^{n+m}$ when $c=1$.
A straightforward computation shows that the standard embedding of a torus
$\mathbb{T}^{n}_{k}(r)=\mathbb{S}^{k}(r)\times\mathbb{S}^{n-k}(\sqrt{1-r^{2}}),\;1\leq k\leq n-1,$
into the unit sphere $\mathbb{S}^{n+1}$, where $\mathbb{S}^{k}(r)$ denotes the $k$-dimensional sphere of radius $r<1$, satisfies ($\ast$1) with equality if $r\geq\sqrt{k/n}$, or $n=2k$. In all other cases, one has $S>a(n,k,H,1)$.
The pinching condition ($\ast$1) has been studied primarily in the case of the smallest admissible value of $k$ (see, e.g., [29, 31, 32, 34]). Shiohama and Xu [25] proved that compact submanifolds in space forms of nonnegative curvature are homeomorphic to a sphere, provided ($\ast$1) holds as a strict inequality at every point when $k=1$. In [27], a homology vanishing result was established for submanifolds in space forms of nonnegative curvature that satisfy the strict pointwise inequality in ($\ast$1) for some integer $1\leq k\leq n-1$. The argument there relied on the nonexistence of stable currents, as established by Lawson and Simons [12], under suitable upper bounds for the second fundamental form.
In our recent work [28], we studied the geometric and topological rigidity of submanifolds of dimension $n\geq 5$ satisfying the pinching condition ($\ast$1), without assuming strictness or imposing any restrictions on the mean curvature. Among other results, we established the following theorem, showing that the pinching condition either forces the vanishing of homology in certain intermediate dimensions or uniquely determines the submanifold up to congruence. Recall that a submanifold is said to be substantial if it is not contained in any proper totally geodesic submanifold of the ambient space.
Theorem 1.
Let $f\colon M^{n}\to\mathbb{Q}_{c}^{n+m}$, $n\geq 5,c\geq 0$, be a substantial isometric immersion of a compact oriented Riemannian manifold. Assume that inequality ($\ast$1) is satisfied for some integer integer $2\leq k\leq n/2$ at every point. Then, either the homology groups of $M^{n}$ satisfy
$H_{p}(M^{n};\mathbb{Z})=0\;\,\text{for all}\;\,k\leq p\leq n-k\;\,\text{and}\;\,H_{k-1}(M^{n};\mathbb{Z})\cong\mathbb{Z}^{\beta_{k-1}(M)},$
where $\beta_{k-1}(M)$ is the (k1)(k-1)-th Betti number, or equality holds in ($\ast$1) at every point and one of the following assertions holds:
$(i)$ $M^{n}$ is isometric to a Clifford torus $\mathbb{T}^{n}_{p}(\sqrt{p/n}),k\leq p\leq n/2$, and $f$ is the standard minimal embedding into $\mathbb{S}^{n+1}$.
$(ii)$ $M^{n}$ is isometric to a torus $\mathbb{T}^{n}_{k}(r)$ with $r>\sqrt{k/n}$ and $f$ is the standard embedding into $\mathbb{S}^{n+1}$.
$(iii)$ $M^{n}$ is isometric to a torus $\mathbb{S}^{k}(r)\times\mathbb{S}^{k}(\sqrt{R^{2}-r^{2}})$ and $f$ is a composition $f=j\circ g$, where $g\colon M^{n}\to\mathbb{S}^{n+1}(R)$ is the standard embedding of the torus $\mathbb{S}^{k}(r)\times\mathbb{S}^{k}(\sqrt{R^{2}-r^{2}})$ into a sphere $\mathbb{S}^{n+1}(R)$ and $j\colon\mathbb{S}^{n+1}(R)\to\mathbb{Q}_{c}^{n+2}$ is an umbilical inclusion with $c=0,1$ and $R<1$ if $c=1$.
The main focus of this paper is the study of four-dimensional submanifolds satisfying condition ($\ast$1), which are not covered by Theorem 1. The results obtained here, as well as the methods employed, differ substantially from those in the proof of Theorem 1 for higher dimensions and are considerably more intricate. Our approach relies essentially on tools from four-dimensional geometry, the theory of Riemannian manifolds with nonnegative isotropic curvature, and the Bochner technique. A key observation is that four-dimensional submanifolds satisfying the pinching condition ($\ast$1) necessarily have nonnegative isotropic curvature, a concept introduced and studied by Micallef and Moore [17].
Throughout the paper, all submanifolds under consideration are assumed to be connected. The main result of the paper can be stated as follows.
Theorem 2.
Let $f\colon M^{4}\to\mathbb{Q}_{c}^{4+m},c\geq 0$, be a substantial isometric immersion of a compact, oriented Riemannian four-manifold. Suppose inequality ($\ast$1) is satisfied for $k=2$ at every point. Then, one of the following assertions holds:
$(i)$ $M^{4}$ is diffeomorphic to $\mathbb{S}^{4}$.
$(ii)$ The universal cover of $M^{4}$ is isometric to a Riemannian product $\mathbb{R}\times N$, where $N$ is diffeomorphic to $\mathbb{S}^{3}$ with nonnegative Ricci curvature.
$(iii)$ Equality holds in ($\ast$1) for $k=2$ at every point, and one of the following holds:
  1. (a)
    $M^{4}$ is isometric to a torus $\mathbb{S}^{2}(r)\times\mathbb{S}^{2}(\sqrt{R^{2}-r^{2}})$ and $f$ is a composition $f=j\circ g$, where $g\colon M^{4}\to\mathbb{S}^{5}(R)$ is the standard embedding of the torus $\mathbb{S}^{2}(r)\times\mathbb{S}^{2}(\sqrt{R^{2}-r^{2}})$ into a sphere $\mathbb{S}^{5}(R)$, and $j\colon\mathbb{S}^{5}(R)\to\mathbb{Q}_{c}^{6}$ is an umbilical inclusion with $R\leq 1$ if $c=1$.
  2. (b)
    $M^{4}$ is isometric to the projective plane $\mathord{\mathbb{C}}\mathord{\mathbb{P}}_{r}^{2}$ of constant holomorphic curvature $4/3r^{2}$ with $r=1/\sqrt{c+H^{2}}$ and $f=j\circ g$, where $g$ is the standard minimal embedding of $\mathord{\mathbb{C}}\mathord{\mathbb{P}}_{r}^{2}$ into $\mathbb{S}^{7}(r)$, and $j\colon\mathbb{S}^{7}(r)\to\mathbb{Q}_{c}^{8}$ is an umbilical inclusion.
In the final section of the paper, we present a method for constructing geometrically distinct isometric immersions of manifolds diffeomorphic either to $\mathbb{S}^{4}$ or to the torus $\mathbb{S}^{1}\times\mathbb{S}^{3}$, which also satisfy ($\ast$1) for $k=2$. These examples demonstrate that Theorem 2 is sharp. In particular, the standard isometric embedding of the torus $\mathbb{S}^{1}(r)\times\mathbb{S}^{3}(\sqrt{1-r^{2}})$ into $\mathbb{S}^{5}$ satisfies ($\ast$1) for $k=2$ whenever $r\geq 1/2$, and its fundamental group is infinite. This example shows that the class of submanifolds appearing in part $(ii)$ of Theorem 2 is nonempty.
For submanifolds of arbitrary dimension satisfying the pinching condition ($\ast$1) with $k=2$, we establish the following result.
Theorem 3.
Let $f\colon M^{n}\to\mathbb{Q}_{c}^{n+m}$, with $n\geq 4$ and $c\geq 0$, be a substantial isometric immersion of a compact, oriented Riemannian manifold. Assume that inequality ($\ast$1) is satisfied at every point for $k=2$. Then, one of the following holds:
$(i)$ The manifold $M^{n}$ is diffeomorphic to $\mathbb{S}^{4}$.
$(ii)$ The universal cover of $M^{n}$ is isometric to a Riemannian product $\mathbb{R}\times N$, where $N$ is diffeomorphic to $\mathbb{S}^{3}$ with nonnegative Ricci curvature.
$(iii)$ If $n\geq 5$, the homology of $M^{n}$ satisfies
$H_{p}(M^{n};\mathbb{Z})=0\;\;\text{for all}\;\;2\leq p\leq n-2$
and
$H_{p}(M^{n};\mathbb{Z})\cong\mathbb{Z}^{\beta_{1}(M)}\;\;\text{for}\;\;p=1,n-1.$
$(iv)$ Equality holds in condition ($\ast$1) for $k=2$ at every point, and the submanifold is as described in parts (i)(i)-$(iii)$ of Theorem 1, or as in part $(iii)$ of Theorem 2.
The results presented here extend those of previous studies [1, 2, 4, 11, 14, 26, 32, 33, 34] without imposing additional assumptions on either the mean curvature or the fundamental group of the submanifold. In the final section, we further construct geometrically distinct isometric immersions of manifolds diffeomorphic to either the sphere $\mathbb{S}^{n}$ or the torus $\mathbb{S}^{1}\times\mathbb{S}^{n-1}$, with $n\geq 4$, which also satisfy ($\ast$1) for $c=0$ and $k=2$. These examples illustrate the sharpness of Theorem 3. Finally, note that the homology described in part $(iii)$ of Theorem 3 coincides with that of the sphere $\mathbb{S}^{n}$ with $g=\beta_{1}(M)$ handles attached, or equivalently, with the homology of the connected sum $_{g}(\mathbb{S}^{1}\times\mathbb{S}^{n-1})$ consisting of $g$ summands.
The geometric and topological rigidity of submanifolds of arbitrary dimension satisfying the pinching condition ($\ast$1) for $k=1$ was studied in [28]. Combined with the results presented here, this provides a comprehensive understanding of compact submanifolds satisfying ($\ast$1) for the two smallest admissible values of the integer $k$.
It is worth noting that in the recent paper [15], the authors extended the pinching result of Simons [26] and that of Chern, do Carmo, and Kobayashi [4] to complete (not necessarily compact) minimal submanifolds in spheres. It would be of interest to extend the results of the present paper, as well as those in [28], to complete submanifolds in space forms, without imposing any restrictions on the mean curvature. Such a generalization could provide deeper insights into the geometry and topology of submanifolds satisfying the pinching condition ($\ast$1), beyond the minimal or compact setting.
2 Background
2.1 Geometry of 4-dimensional manifolds
In this section, we collect basic facts about four-dimensional geometry. For a detailed exposition of the subject, we refer the reader to [13, 24].
Let $(M,{\langle}\cdot,\cdot{\rangle})$ be an oriented Riemannian manifold of dimension $n=4$ with Levi-Civitá connection $\nabla$ and curvature tensor $R$ given by
$R(X,Y)=[\nabla_{X},\nabla_{Y}]-\nabla_{[X,Y]},\;X,Y\in\mathcal{X}(M).$
The Ricci tensor of $(M,{\langle}\cdot,\cdot{\rangle})$ is defined by
$\mbox{Ric}(X,Y)=\sum_{i}{\langle}R(X,E_{i})E_{i},Y{\rangle},\;X,Y\in\mathcal{X}(M),$
where $\{E_{i}\}_{1\leq i\leq 4}$ is a local orthonormal frame.
At any point $x\in M$, we consider the Bochner-Weitzenböck operator $\mathcal{B}^{[2]}$ as an endomorphism of the space of 2-vectors $\Lambda^{2}T_{x}M$ at $x$ given by
${\langle}{\langle}\mathcal{B}^{[2]}(v_{1}\wedge v_{2}),w_{1}\wedge w_{2}{\rangle}{\rangle}=$ $\;\mbox{Ric}(v_{1},w_{1}){\langle}v_{2},w_{2}{\rangle}+\mbox{Ric}(v_{2},w_{2}){\langle}v_{1},w_{1}{\rangle}$
$-\mbox{Ric}(v_{1},w_{2}){\langle}v_{2},w_{1}{\rangle}-\mbox{Ric}(v_{2},w_{1}){\langle}v_{1},w_{2}{\rangle}$
$-2{\langle}R(v_{1},v_{2})w_{2},w_{1}{\rangle},$ (1)
and then extend it linearly to all of $\Lambda^{2}T_{x}M$. Here $\langle$$\langle$$\cdot$$\cdot{\rangle}{\rangle}$ stands for the inner product of $\Lambda^{2}T_{x}M$ defined by
${\langle}{\langle}v_{1}\wedge v_{2},w_{1}\wedge w_{2}{\rangle}{\rangle}=\det({\langle}v_{i},w_{j}{\rangle}).$
The Bochner-Weitzenböck operator is a self-adjoint operator. If $Z\in\Lambda^{2}T_{x}M$, the dual 2-form $\omega$ is defined by
$\omega(v,w)={\langle}{\langle}Z,v\wedge w{\rangle}{\rangle},$
and we may regard $Z$ as the skew-symmetric endomorphism of the tangent space at $x$ via
${\langle}Z(v),w{\rangle}={\langle}{\langle}Z,v\wedge w{\rangle}{\rangle}.$
Clearly $\mathcal{B}^{[2]}$ can also be viewed as an endomorphism of the bundle $\Omega^{2}(M)$ of 2-forms of the manifold via the inner product $\langle$$\langle$$\cdot$$\cdot{\rangle}{\rangle}$. If $\omega$ is a 2-form, then $\mathcal{B}^{[2]}(\omega)$ is given by
$\mathcal{B}^{[2]}(\omega)(X_{1},X_{2})=\omega(\mbox{Ric}(X_{1}),X_{2})+\omega(X_{1},\mbox{Ric}(X_{2}))-\sum_{i}\omega(R(X_{1},X_{2})E_{i},E_{i}).$
Then the Bochner-Weitzenböck operator acts on $\Lambda^{2}TM$ by
$\omega(\mathcal{B}^{[2]}(X_{1}\wedge X_{2}))=\mathcal{B}^{[2]}(\omega)(X_{1},X_{2}).$
Taking $\omega$ to be the dual form to $w_{1}\wedge w_{2}$ yields (2.1). Throughout the paper, we will generally identify 2-forms with 2-vectors.
We recall the Bochner-Weitzenböck formula which can be written as
${\langle}\Delta\omega,\omega{\rangle}=\frac{1}{2}\Delta\|\omega\|^{2}+\|\nabla\omega\|^{2}+{\langle}\mathcal{B}^{[2]}(\omega),\omega{\rangle}$
for any $\omega\in\Omega^{2}(M)$. This implies that any harmonic 2-form on a compact manifold is parallel provided that the Bochner-Weitzenböck operator is nonnegative.
The bundle of 2-forms of any oriented 4-dimensional Riemannian manifold $M$ decomposes as a direct sum
$\Omega^{2}(M)=\Omega_{+}^{2}(M)\oplus\Omega_{-}^{2}(M)$
of the eigenspaces of the Hodge star operator
$\ast\colon\Omega^{2}(M)\to\Omega^{2}(M)$
The sections of $\Omega_{+}^{2}(M)$ are called self-dual 2-forms, whereas the ones of $\Omega_{-}^{2}(M)$ are called anti-self-dual 2-forms. Accordingly, at any point $x\in M$, we have the splitting
$\Lambda^{2}T_{x}M=\Lambda_{+}^{2}T_{x}M\oplus\Lambda_{-}^{2}T_{x}M,$
where $\Lambda_{\pm}^{2}T_{x}M$ are the eigenspaces of the Hodge star operator
$\ast\colon\Lambda^{2}T_{x}M\to\Lambda^{2}T_{x}M.$
Both spaces $\Lambda_{+}^{2}T_{x}M$ and $\Lambda_{-}^{2}T_{x}M$ are $\mathcal{B}^{[2]}$-invariant (see Proposition 1 in [23]). Hence, we obtain the corresponding decomposition
$\mathcal{B}^{[2]}=\mathcal{B}_{+}^{[2]}\oplus\mathcal{B}_{-}^{[2]}.$
Suppose now that $M$ is a compact, oriented Riemannian four-manifold. The Hodge theorem guarantees that every de Rham cohomology class on $M$ has a unique harmonic representative. In particular, the space $\mathscr{H}^{2}(M)$ of harmonic 2-forms decomposes as
$\mathscr{H}^{2}(M)=\mathscr{H}^{2}_{+}(M)\oplus\mathscr{H}^{2}_{-}(M),$
where $\mathscr{H}^{2}_{+}(M)$ and $\mathscr{H}^{2}_{-}(M)$ are the spaces of self-dual and anti-self-dual harmonic 2-forms, respectively. The dimensions of these subspaces, denoted by $\beta_{\pm}(M)=\dim\mathscr{H}^{2}_{\pm}(M)$, are oriented homotopy invariants of $M$. Their difference
$\sigma=\beta_{+}(M)-\beta_{-}(M)$
is the signature of $M$, while their sum equals the second Betti number $\beta_{2}(M)$ of the manifold.
2.2 Isotropic curvature
Let $(M,g)$ be a Riemannian manifold of dimension $n\geq 4$. We say that $(M,g)$ has nonnegative isotropic curvature at a point $x\in M$ if
$R_{1331}+R_{1441}+R_{2332}+R_{2442}-2R_{1234}\geq 0,$
for all orthonormal four-frames $\{e_{1},e_{2},e_{3},e_{4}\}\subset T_{x}M$. Here, $R_{ijk\ell}$ are the components of the curvature tensor $R$, defined by
$R_{ijk\ell}=g\big{(}R(e_{i},e_{j})e_{k},e_{\ell}\big{)},\;1\leq i,j,k,\ell\leq 4.$
If the strict inequality holds, we say that $(M,g)$ has positive isotropic curvature at $x$. The manifold $(M,g)$ is said to have nonnegative (or positive) isotropic curvature if it satisfies the corresponding condition at every point and for all orthonormal four-frames.
The following result is well known (see, for instance, [19]).
Lemma 4.
For any four-dimensional Riemannian manifold $M$, the non-negativity of the isotropic curvature at a point $x\in M$ is equivalent to the non-negativity of the Bochner-Weitzenböck operator $\mathcal{B}^{[2]}$ at $x$.
2.3 The pinching condition and the Lawson-Simons inequality
Lawson and Simons [12] proved that specific bounds on the second fundamental form for submanifolds of spheres can result to the vanishing of homology groups with integer coefficients. By employing the second variation of area, they effectively ruled out stable minimal currents in certain dimensions. Consequently, since the existences theorems of Federer and Fleming [8] allow area minimization within a homology class, this result implies the triviality of integral homology.
The result of Lawson and Simons [12] mentioned above was later strengthened by Elworthy and Rosenberg [6, p. 71] without requiring the bound on the second fundamental form to be strict at all points of the submanifold. In this section, we state their theorem and then quote an auxiliary result from [28] that establishes the relation between our pinching condition ($\ast$1) and the inequality ($\ast\ast$5) below required in their result.
Theorem 5.
([6, 12, 30]) Let $f\colon M^{n}\to\mathbb{Q}_{c}^{n+m},n\geq 4,c\geq 0$, be an isometric immersion of a compact Riemannian manifold. Let $p$ be an integer such that $1\leq p\leq n-1$. Assume that at any point $x\in M^{n}$ and for every orthonormal basis $\{e_{1},\ldots,e_{n}\}$ of $T_{x}M$, the second fundamental form $\alpha_{f}\colon TM\times TM\to N_{f}M$ satisfies
$\sum_{i=1}^{p}\sum_{j=p+1}^{n}\big{(}2\|\alpha_{f}(e_{i},e_{j})\|^{2}-{\langle}\alpha_{f}(e_{i},e_{i}),\alpha_{f}(e_{j},e_{j}){\rangle}\big{)}\leq p(n-p)c.$ ($\ast\ast$)
If there exists a point at which inequality ($\ast\ast$5) is strict for every orthonormal basis of the tangent space at that point, then the homology groups satisfy
$H_{p}(M^{n};\mathbb{Z})=H_{n-p}(M^{n};\mathbb{Z})=0.$
Next, we recall the following result from [28], which establishes the relationship between the inequalities ($\ast$1) and ($\ast\ast$5) for four-dimensional submanifolds.
Lemma 6.
Let $f\colon M^{4}\to\mathbb{Q}_{c}^{4+m},c\geq 0$, be an isometric immersion of a $4$-dimensional Riemannian manifold $M^{4}$ such that inequality ($\ast$1) is satisfied at a point $x\in M^{4}$ for $k=2$. Then the following assertions hold at $x$:
$(i)$ The inequality ($\ast\ast$5) holds for $p=2$ and for all orthonormal bases of $T_{x}M$. Moreover, if inequality ($\ast$1) is strict at $x$, then ($\ast\ast$5) is also strict for every orthonormal basis of $T_{x}M$.
$(ii)$ Suppose now that equality holds in ($\ast\ast$5) for some orthonormal basis $\{e_{i}\}_{1\leq i\leq 4}$ of $T_{x}M$ and $p=2$. Then there exist normal vectors $\eta_{1},\eta_{2}\in N_{f}M(x)$ such that the shape operator $A_{\xi}$ associated to any $\xi\in N_{f}M(x)$ satisfies
$\pi_{j}\circ A_{\xi}|_{V_{j}}={\langle}\xi,\eta_{j}{\rangle}Id,\;j=1,2,$
where $Id$ is the identity map on the tangent space at $x$, $V_{1}=\mbox{span}\{e_{1},e_{2}\}$, $V_{2}=\mbox{span}\{e_{3},e_{4}\}$ and $\pi_{j}$ denotes the projection onto $V_{j},j=1,2$.
3 Proofs of the main results
We recall a key result, namely Proposition 16 from [20], which provides an estimate for the Bochner-Weitzenböck operator of a submanifold in terms of its second fundamental form. For four-dimensional submanifolds, this proposition can be stated as follows:
Proposition 7.
Let $f\colon M^{4}\to\mathbb{Q}_{c}^{4+m}$ be an isometric immersion of a $4$-dimensional manifold $M^{4}$. The Bochner-Weitzenböck operator $\mathcal{B}^{[2]}$ of $M^{4}$ satisfies the following pointwise inequality
$\mathop{\min_{\omega\in\Omega^{2}(M^{4})}}_{\|\omega\|=1}{\langle}\mathcal{B}^{[2]}\omega,\omega{\rangle}\geq 4c+8H^{2}-S.$ (2)
If equality holds in (2) at a point $x\in M^{4}$, then for every unit vector $\xi\in N_{f}M(x)$, the shape operator $A_{\xi}(x)$ has at most two distinct eigenvalues, each of multiplicity $2$.
We now state the following auxiliary results.
Proposition 8.
Let $f\colon M^{4}\to\mathbb{Q}^{4+m}_{c},c\geq 0$, be an isometric immersion of an oriented $4$-dimensional Riemannian manifold. Suppose that $f$ satisfies inequality ($\ast$1) at every point for $k=2$. Then, the following assertions hold:
$(i)$ The Bochner-Weitzenböck operator $\mathcal{B}^{[2]}$ is nonnegative, and the manifold $M^{4}$ consequently has nonnegative isotropic curvature.
$(ii)$ Suppose that $M^{4}$ is compact and satisfies $\beta_{2}(M)\neq 0$. Then equality holds in ($\ast$1) at every point, and the shape operator, associated with any normal direction has at most two distinct eigenvalues, each of multiplicity $2$. Moreover, at each point $x\in M^{4}$, there exists an oriented orthonormal four-frame $\{e_{1},e_{2},e_{3},e_{4}\}\subset T_{x}M$ such that the second fundamental form $\alpha_{f}$ of $f$ satisfies the following conditions:
$\;\;\;\;\alpha_{11}=\alpha_{22},\,\alpha_{33}=$ $\alpha_{44},\,\alpha_{12}=\alpha_{34}=0,\,\|\alpha_{23}\|=\|\alpha_{14}\|,\,\|\alpha_{24}\|=\|\alpha_{13}\|,$ (3)
${\langle}\alpha_{14}+\alpha_{23}$ $,\alpha_{13}-\alpha_{24}{\rangle}=0,\;\,{\langle}\alpha_{13}+\alpha_{24},\alpha_{14}-\alpha_{23}{\rangle}=0,$ (4)
$\|\alpha_{13}\|^{2}+\|\alpha_{14}\|^{2}=c+{\langle}\alpha_{11},\alpha_{44}{\rangle},$ (5)
where, for simplicity, we denote $\alpha_{ij}=\alpha_{f}(e_{i},e_{j})$.
Proof.
$(i)$ From ($\ast$1), if follows that the right-hand side of inequality (2) in Proposition 7 is nonnegative. Consequently, the Bochner-Weitzenböck operator $\mathcal{B}^{[2]}$ is nonnegative. By Lemma 4, this implies that the manifold $M^{4}$ has nonnegative isotropic curvature.
$(ii)$ By the Hodge theorem, there exists a nontrivial harmonic 2-form $\omega\in\Omega^{2}(M)$. The Bochner-Weitzenböck formula then implies
${\langle}\mathcal{B}^{[2]}\omega,\omega{\rangle}=0$
at every point. Therefore, equality holds in inequality (2), and as a result, Proposition 7 implies that the shape operator, associated with any normal direction has at most two distinct eigenvalues, each of multiplicity $2$.
On the other hand, from part $(i)$ of Lemma 6, we know that inequality ($\ast\ast$5) holds for $p=2$ at every point and for every orthonormal basis. Since $H_{2}(M;\mathbb{Z})\neq 0$, Theorem 5 implies that at each point $x\in M^{4}$ there exists an oriented orthonormal four-frame $\{e_{1},e_{2},e_{3},e_{4}\}\subset T_{x}M$ such that equality holds in ($\ast\ast$5) for $p=2$. It follows from part $(i)$ of Lemma 6 that equality also holds in ($\ast$1) at any point. Furthermore, by choosing an orthonormal normal basis $\{\xi_{\alpha}\}_{1\leq\alpha\leq m}$ at $x\in M^{4}$ such that the mean curvature vector is $\mathcal{H}(x)=H\xi_{1}$, part $(ii)$ of Lemma 6 implies that the corresponding shape operators $A_{\xi_{\alpha}},1\leq\alpha\leq m$, are given by
$\begin{cases}A_{\xi_{\alpha}}e_{1}=\rho_{\alpha}e_{1}+\kappa_{\alpha}e_{3}+\lambda_{\alpha}e_{4},\\ A_{\xi_{\alpha}}e_{2}=\rho_{\alpha}e_{2}+\mu_{\alpha}e_{3}+\nu_{\alpha}e_{4},\\ A_{\xi_{\alpha}}e_{3}=\kappa_{\alpha}e_{1}+\mu_{\alpha}e_{2}+\sigma_{\alpha}e_{3},\\ A_{\xi_{\alpha}}e_{4}=\lambda_{\alpha}e_{1}+\nu_{\alpha}e_{2}+\sigma_{\alpha}e_{4},\end{cases}$ (6)
where
$\rho_{1}+\sigma_{1}=2H\;\;\text{and}\;\;\rho_{\alpha}+\sigma_{\alpha}=0\;\;\text{for any}\;\;2\leq\alpha\leq m.$
Since each shape operator has at most two distinct eigenvalues, each of multiplicity $2$, it follows from (6) that
$\nu_{\alpha}=\pm\kappa_{\alpha}\;\;\text{and}\;\;\mu_{\alpha}=\mp\lambda_{\alpha}\;\;\text{for any}\;\;1\leq\alpha\leq m.$
From this, (3) follows directly, along with the relations:
${\langle}\alpha_{13},\alpha_{14}{\rangle}+{\langle}\alpha_{23},\alpha_{24}{\rangle}=0,\;\,{\langle}\alpha_{13},\alpha_{23}{\rangle}+{\langle}\alpha_{14},\alpha_{24}{\rangle}=0.$ (7)
Using (3), (7), and the relation $\alpha_{11}+\alpha_{44}=\mathcal{H}/2$, we deduce that ($\ast$1), which now holds as an equality, is equivalent to (5). Furthermore, if (4) is satisfied, the orthonormal four-frame $\{e_{i}\}_{1\leq i\leq 4}$ satisfies all the desired properties.
Suppose now that the frame $\{e_{i}\}_{1\leq i\leq 4}$ does not satisfy (4). Consider instead the orthonormal four-frame $\{\tilde{e}_{i}\}_{1\leq i\leq 4}$, defined by
$\tilde{e}_{i}=R_{\theta}e_{i},\;\tilde{e}_{j}=R_{\varphi}e_{j},\;i=1,2,\;j=3,4.$
Here, $R_{\theta}$ and $R_{\varphi}$ denote rotations on $V_{1}=\mbox{span}\{e_{1},e_{2}\}$ and $V_{2}=\mbox{span}\{e_{3},e_{4}\}$ by angles $\theta$ and $\varphi$, respectively. For simplicity, we set $\tilde{\alpha}_{ij}=\alpha_{f}(\tilde{e}_{i},\tilde{e}_{j})$. Then, using (3), we have
$\tilde{\alpha}_{ii}=\alpha_{ii},\;1\leq i\leq 4,\;\;{\text{and}}\;\;\tilde{\alpha}_{12}=\tilde{\alpha}_{34}=0.$
Since each shape operator has at most two distinct eigenvalues, each of multiplicity $2$, a similar argument as above shows that the vectors $\tilde{\alpha}_{ij}$, for $1\leq i,j\leq 4$, satisfy (3), (5) and (7). We now claim that the angles $\theta$ and $\varphi$ can be chosen such that (4) is satisfied for the frame $\{\tilde{e}_{i}\}_{1\leq i\leq 4}$. Straightforward computations then yield the following relations:
$\tilde{\alpha}_{14}+\tilde{\alpha}_{23}=\cos(\varphi+\theta)(\alpha_{14}+\alpha_{23})+\sin(\varphi+\theta)(\alpha_{24}-\alpha_{13}),$ (8)
$\tilde{\alpha}_{24}-\tilde{\alpha}_{13}=-\sin(\varphi+\theta)(\alpha_{14}+\alpha_{23})+\cos(\varphi+\theta)(\alpha_{24}-\alpha_{13}),$ (9)
$\tilde{\alpha}_{13}+\tilde{\alpha}_{24}=\cos(\varphi-\theta)(\alpha_{13}+\alpha_{24})+\sin(\varphi-\theta)(\alpha_{14}-\alpha_{23}),$ (10)
$\tilde{\alpha}_{14}-\tilde{\alpha}_{23}=-\sin(\varphi-\theta)(\alpha_{13}+\alpha_{24})+\cos(\varphi-\theta)(\alpha_{14}-\alpha_{23}).$ (11)
We can select angles $\sigma_{1}$ and $\sigma_{2}$ such that the following conditions are satisfied:
$2\cos\sigma_{1}{\langle}\alpha_{14}+\alpha_{23},\alpha_{24}-\alpha_{13}{\rangle}+\sin\sigma_{2}\big{(}\|\alpha_{13}-\alpha_{24}\|^{2}-\|\alpha_{14}+\alpha_{23}\|^{2}\big{)}=0,$
$2\cos\sigma_{2}{\langle}\alpha_{13}+\alpha_{24},\alpha_{14}-\alpha_{23}{\rangle}-\sin\sigma_{2}\big{(}\|\alpha_{14}-\alpha_{23}\|^{2}-\|\alpha_{13}+\alpha_{24}\|^{2}\big{)}=0.$
By choosing
$\varphi=\frac{1}{4}(\sigma_{1}+\sigma_{2})\;\;{\text{and}}\;\;\theta=\frac{1}{4}(\sigma_{1}-\sigma_{2}),$
and applying (8)-(11), it follows that
${\langle}\tilde{\alpha}_{14}+\tilde{\alpha}_{23},\tilde{\alpha}_{13}-\tilde{\alpha}_{14}{\rangle}=0,\;\,{\langle}\tilde{\alpha}_{13}+\tilde{\alpha}_{24},\tilde{\alpha}_{14}-\tilde{\alpha}_{23}{\rangle}=0.$
This completes the proof of the proposition.∎
We recall that the first normal space of an isometric immersion $f\colon M^{n}\to\mathbb{Q}_{c}^{n+m}$ at a point $x\in M^{n}$ is defined by
$N_{1}f(x)=\operatorname{span}\left\{\alpha_{f}(X,Y):X,Y\in T_{x}M\right\}.$
Lemma 9.
Let $f\colon M^{4}\to\mathbb{Q}^{4+m}_{c},c\geq 0$, be an isometric immersion of an oriented $4$-dimensional Riemannian manifold $M^{4}$. Suppose that at a point $x\in M^{4}$, there exists an oriented orthonormal four-frame $\{e_{1},e_{2},e_{3},e_{4}\}\subset T_{x}M$ such that the second fundamental form of $f$ at $x$ satisfies conditions (3)-(5) in part (ii) of Proposition 8. We consider the orthonormal basis $\{\eta_{i}\}_{1\leq i\leq 6}$ of the space of 2-vectors $\Lambda^{2}T_{x}M$, defined by the relations
$\eta_{i}\in\Lambda_{+}^{2}T_{x}M,\;\;\eta_{i+3}=\ast\eta_{i},\;\;\text{for}\;\;1\leq i\leq 3,$
and
$\eta_{1}=\frac{1}{\sqrt{2}}(e_{12}+e_{34}),\;\eta_{2}$ $=\frac{1}{\sqrt{2}}(e_{13}-e_{24}),\;\eta_{3}=\frac{1}{\sqrt{2}}(e_{14}+e_{23}),$
where $e_{ij}=e_{i}\wedge e_{j}$. Then the following assertions hold at $x$:
$(i)$ The matrix of the Bochner-Weitzenböck operator $\mathcal{B}^{[2]}=\mathcal{B}_{+}^{[2]}\oplus\mathcal{B}_{-}^{[2]}$ at the point $x$, with respect to the basis $\{\eta_{i}\}_{1\leq i\leq 6}$, is given by the direct sum
$\begin{bmatrix}\mu^{+}_{1}&a^{+}_{1}&a^{+}_{2}\\ a^{+}_{1}&\mu^{+}_{2}&0\\ a^{+}_{2}&0&\mu^{+}_{3}&\!\!\!\!\!\end{bmatrix}\oplus\begin{bmatrix}\mu^{-}_{1}&a^{-}_{1}&a^{-}_{2}\\ a^{-}_{1}&\mu^{-}_{2}&0\\ a^{-}_{2}&0&\mu^{-}_{3}&\!\!\!\!\!\end{bmatrix}$,
where
$\mu^{\pm}_{1}=\;$ $\|\alpha_{14}\pm\alpha_{23}\|^{2}+\|\alpha_{13}\mp\alpha_{24}\|^{2},$
$\mu^{\pm}_{2}=\;$ $\|\alpha_{11}-\alpha_{44}\|^{2}+4\|\alpha_{13}\|^{2}+2\left(\|\alpha_{14}\|\|\alpha_{23}\|\mp{\langle}\alpha_{14},\alpha_{23}{\rangle}\right),$
$\mu^{\pm}_{3}=\;$ $\|\alpha_{11}-\alpha_{44}\|^{2}+4\|\alpha_{14}\|^{2}+2\left(\|\alpha_{13}\|\|\alpha_{24}\|\pm{\langle}\alpha_{13},\alpha_{24}{\rangle}\right),$
$a^{\pm}_{1}=\;$ ${\langle}\alpha_{23}\pm\alpha_{14},\alpha_{44}-\alpha_{11}{\rangle},\;\,a^{\pm}_{2}={\langle}\alpha_{24}\mp\alpha_{13},\alpha_{44}-\alpha_{11}{\rangle}.$
$(ii)$ If $\ker\mathcal{B}_{+}^{[2]}\neq 0$ at $x$, then one of the following conditions holds:
$a_{14}+\alpha_{23}=0=\alpha_{13}-\alpha_{24}\;\;{\text{a}nd}\;\;a_{1}^{+}=a_{2}^{+}=0,$ (ii1)
$\alpha_{13}=\alpha_{24}=0,\;\alpha_{23}=\alpha_{14}\neq 0\;\;{\text{a}nd}\;\;\alpha_{44}-\alpha_{11}=2\rho\,\alpha_{14},\;\rho\in\mathbb{R},$ (ii2)
$\alpha_{14}=\alpha_{23}=0,\;\alpha_{24}=-\alpha_{13}\neq 0\;\;{\text{a}nd}\;\;\alpha_{44}-\alpha_{11}=2\rho\,\alpha_{13},\;\rho\in\mathbb{R}.$ (ii3)
$(iii)$ If $\ker\mathcal{B}_{-}^{[2]}\neq 0$ at $x$, then one of the following conditions holds:
$\alpha_{14}-\alpha_{23}=0=\alpha_{13}+\alpha_{24}\;\;{\text{a}nd}\;\;a_{1}^{-}=a_{2}^{-}=0,$ (iii1)
$\alpha_{13}=\alpha_{24}=0,\;\alpha_{23}=-\alpha_{14}\neq 0\;\;{\text{a}nd}\;\;a_{44}-\alpha_{11}=2\rho\,\alpha_{14},\;\rho\in\mathbb{R},$ (iii2)
$\alpha_{14}=\alpha_{23}=0,\;\alpha_{24}=\alpha_{23}\neq 0\;\;{\text{a}nd}\;\;\alpha_{44}-\alpha_{11}=2\rho\,\alpha_{13},\;\rho\in\mathbb{R}.$ (iii3)
In addition, in cases (ii2), (ii3), (iii2), and (iii3), the immersion $f$ has flat normal bundle at $x$, and its first normal space satisfies
$1\leq\dim N_{1}f(x)\leq 2.$
Proof.
$(i)$ By a straightforward computation using (2.1), the Gauss equation, and (3)-(5), we obtain the following:
$\mathcal{B}^{[2]}(\eta_{1})=$  $\;2\big{(}2c+2{\langle}\alpha_{11},\alpha_{44}{\rangle}-\|\alpha_{13}\|^{2}-\|\alpha_{14}\|^{2}+{\langle}\alpha_{14},\alpha_{23}{\rangle}-{\langle}\alpha_{13},\alpha_{24}{\rangle}\big{)}\eta_{1}$
$+{\langle}\alpha_{14}+\alpha_{23},\alpha_{44}-\alpha_{11}{\rangle}\eta_{2}+{\langle}\alpha_{24}-\alpha_{13},\alpha_{44}-\alpha_{11}{\rangle}\eta_{3}$
$=$ $\;\big{(}\|\alpha_{14}+\alpha_{23}\|^{2}+\|\alpha_{13}-\alpha_{24}\|^{2}\big{)}\eta_{1}+a_{1}^{+}\eta_{2}+a_{2}^{+}\eta_{3},$
$\mathcal{B}^{[2]}(\eta_{2})=$ $\;{\langle}\alpha_{14}+\alpha_{23},\alpha_{44}-\alpha_{11}{\rangle}\eta_{1}$
$+\big{(}4c+\|\alpha_{11}\|^{2}+\|\alpha_{44}\|^{2}+2{\langle}\alpha_{11},\alpha_{44}{\rangle}-2\|\alpha_{14}\|^{2}-2{\langle}\alpha_{14},\alpha_{23}{\rangle}\big{)}\eta_{2}$
$=$ $\;a_{1}^{+}\eta_{1}+\big{(}\|\alpha_{11}-\alpha_{44}\|^{2}+4\|\alpha_{13}\|^{2}+2\|\alpha_{14}\|\|\alpha_{23}\|-2{\langle}\alpha_{14},\alpha_{23}{\rangle}\big{)}\eta_{2},$
$\mathcal{B}^{[2]}(\eta_{3})=$ $\;{\langle}\alpha_{24}-\alpha_{13},\alpha_{44}-\alpha_{11}{\rangle}\eta_{1}$
$+\big{(}4c+\|\alpha_{11}\|^{2}+\|\alpha_{44}\|^{2}+2{\langle}\alpha_{11},\alpha_{44}{\rangle}-2\|\alpha_{13}\|^{2}+2{\langle}\alpha_{13},\alpha_{24}{\rangle}\big{)}\eta_{3}$
$=$ $\;a_{2}^{+}\eta_{1}+\big{(}\|\alpha_{11}-\alpha_{44}\|^{2}+4\|\alpha_{14}\|^{2}+2\|\alpha_{13}\|\|\alpha_{24}\|+2{\langle}\alpha_{13},\alpha_{24}{\rangle}\big{)}\eta_{3},$
and
$\mathcal{B}^{[2]}(\eta_{4})=$  $\;2\big{(}2c+2{\langle}\alpha_{11},\alpha_{44}{\rangle}-\|\alpha_{13}\|^{2}-\|\alpha_{14}\|^{2}-{\langle}\alpha_{14},\alpha_{23}{\rangle}+{\langle}\alpha_{13},\alpha_{24}{\rangle}\big{)}\eta_{4}$
$+{\langle}\alpha_{23}-\alpha_{14},\alpha_{44}-\alpha_{11}{\rangle}\eta_{5}+{\langle}\alpha_{24}+\alpha_{13},\alpha_{44}-\alpha_{11}{\rangle}\eta_{6}$
$=$ $\;\big{(}\|\alpha_{14}-\alpha_{23}\|^{2}+\|\alpha_{13}+\alpha_{24}\|^{2}\big{)}\eta_{4}+a_{1}^{-}\eta_{5}+a_{2}^{-}\eta_{6},$
$\mathcal{B}^{[2]}(\eta_{5})=$ $\;{\langle}\alpha_{23}-\alpha_{14},\alpha_{44}-\alpha_{11}{\rangle}\eta_{4}$
$+\big{(}4c+\|\alpha_{11}\|^{2}+\|\alpha_{44}\|^{2}+2{\langle}\alpha_{11},\alpha_{44}{\rangle}-2\|\alpha_{14}\|^{2}+2{\langle}\alpha_{14},\alpha_{23}{\rangle}\big{)}\eta_{5}$
$=$ $\;a_{1}^{-}\eta_{4}+\big{(}\|\alpha_{11}-\alpha_{44}\|^{2}+4\|\alpha_{13}\|^{2}+2\|\alpha_{14}\|\|\alpha_{23}\|+2{\langle}\alpha_{14},\alpha_{23}{\rangle}\big{)}\eta_{5},$
$\mathcal{B}^{[2]}(\eta_{6})=$ $\;{\langle}\alpha_{24}+\alpha_{13},\alpha_{44}-\alpha_{11}{\rangle}\eta_{4}$
$+\big{(}4c+\|\alpha_{11}\|^{2}+\|\alpha_{44}\|^{2}+2{\langle}\alpha_{11},\alpha_{44}{\rangle}-2\|\alpha_{13}\|^{2}-2{\langle}\alpha_{13},\alpha_{24}{\rangle}\big{)}\eta_{6}$
$=$ $\;a_{2}^{-}\eta_{4}+\big{(}\|\alpha_{11}-\alpha_{44}\|^{2}+4\|\alpha_{14}\|^{2}+2\|\alpha_{13}\|\|\alpha_{24}\|-2{\langle}\alpha_{13},\alpha_{24}{\rangle}\big{)}\eta_{6}.$
This completes the proof of part $(i)$.
$(ii)$ Using part $(i)$, we find that
$\det\mathcal{B}_{+}^{[2]}=\mu^{+}_{1}\mu^{+}_{2}\mu^{+}_{3}-\mu^{+}_{2}(a_{2}^{+})^{2}-\mu^{+}_{3}(a_{1}^{+})^{2}.$
Hence, if $\ker\mathcal{B}_{+}^{[2]}\neq 0$ at $x$, then
$\mu^{+}_{1}\mu^{+}_{2}\mu^{+}_{3}=\mu^{+}_{2}(a_{2}^{+})^{2}+\mu^{+}_{3}(a_{1}^{+})^{2},$
or equivalently
$\mu^{+}_{2}\big{(}\mu^{+}_{3}\|\alpha_{13}-\alpha_{24}\|^{2}-(a_{2}^{+})^{2}\big{)}+\mu^{+}_{3}\big{(}\mu^{+}_{2}\|\alpha_{14}+\alpha_{23}\|^{2}-(a_{1}^{+})^{2}\big{)}=0.$ (12)
On the other hand, by applying the Cauchy-Schwarz inequality, we obtain
$\mu^{+}_{2}\big{(}\mu^{+}_{3}\|\alpha_{13}-\alpha_{24}\|^{2}$ $-(a_{2}^{+})^{2}\big{)}+\mu^{+}_{3}\big{(}\mu^{+}_{2}\|\alpha_{14}+\alpha_{23}\|^{2}-(a_{1}^{+})^{2}\big{)}$
$\geq$ $\;\mu^{+}_{2}\|\alpha_{13}-\alpha_{24}\|^{2}\big{(}\mu^{+}_{3}-\|\alpha_{44}-\alpha_{11}\|^{2}\big{)}$
$+\mu^{+}_{3}\|\alpha_{14}+\alpha_{23}\|^{2}\big{(}\mu^{+}_{2}-\|\alpha_{44}-\alpha_{11}\|^{2}\big{)}$
$=$  $\;2\mu^{+}_{2}\|\alpha_{13}-\alpha_{24}\|^{2}\big{(}2\|\alpha_{14}\|^{2}+\|\alpha_{13}\|\|\alpha_{24}\|+{\langle}\alpha_{13},\alpha_{24}{\rangle}\big{)}$
$+2\mu^{+}_{3}\|\alpha_{14}+\alpha_{23}\|^{2}\big{(}2\|\alpha_{13}\|^{2}+\|\alpha_{14}\|\|\alpha_{23}\|-{\langle}\alpha_{14},\alpha_{23}{\rangle}\big{)}.$
The above, together with (12), implies that
$\mu^{+}_{2}\|\alpha_{13}-\alpha_{24}\|^{2}\big{(}2\|\alpha_{14}\|^{2}+\|\alpha_{13}\|\|\alpha_{24}\|+{\langle}\alpha_{13},\alpha_{24}{\rangle}\big{)}=0,$
$\mu^{+}_{3}\|\alpha_{14}+\alpha_{23}\|^{2}\big{(}2\|\alpha_{13}\|^{2}+\|\alpha_{14}\|\|\alpha_{23}\|-{\langle}\alpha_{14},\alpha_{23}{\rangle}\big{)}=0.$
This clearly concludes the proof of part $(ii)$. The proof of part $(iii)$ proceeds similarly and is therefore omitted.
The Ricci equation states that the normal curvature tensor $R^{\perp}$ of the submanifold $f$ is given by
$R^{\perp}(X,Y)\xi=\alpha_{f}(X,A_{\xi}Y)-\alpha_{f}(A_{\xi}X,Y)$
for all $X,Y\in\mathcal{X}(M)$ and $\xi\in\Gamma(N_{f}M)$. We claim that
$\alpha_{f}(e_{i},A_{\xi}e_{j})-\alpha_{f}(A_{\xi}e_{i},e_{j})=0\;\;\text{for all}\;\;1\leq i,j\leq 4,$ (13)
which clearly implies that the immersionff has flat normal bundle at the point $x$.
We prove (13) in the case where (ii2) holds, since the same argument applies to the other cases. Let $(h_{ij})$ denote the matrix of the shape operator $A_{\xi}$ with respect to the four-frame $\{e_{1},e_{2},e_{3},e_{4}\}\subset T_{x}M$. Then it follows from part (ii2) of Lemma 9 that
$h_{13}=h_{24}=0,\;h_{23}=-h_{14}\neq 0\;\;\text{and}\;\;h_{44}-h_{11}=2\rho\,h_{14}.$
That (13) holds follows directly from these relations together with part (ii2) of Lemma 9.
Finally, the inequality on the dimension of the first normal space at $x$ follows directly from parts (ii2), (ii3), (iii2) and (iii3) of Lemma 9. ∎
Let $f\colon M^{n}\to\mathbb{Q}_{c}^{n+m}$ be an isometric immersion. We recall that a vector $\eta$ in the normal space $N_{f}M(x)$ is called a Dupin principal normal of the isometric immersion $f$ at a point $x\in M^{n}$ if the tangent subspace
$E_{\eta}(x)=\left\{X\in T_{x}M:\alpha_{f}(X,Y)={\langle}X,Y{\rangle}\eta\;\;\text{for all}\;\,Y\in T_{x}M\right\}$
is at least two dimensional. That dimension is called the multiplicity of $\eta$. The relative nullity subspace $\mathcal{D}_{f}(x)$ of $f$ at a point $x\in M^{n}$ is the kernel of the second fundamental form at this point, namely
$\mathcal{D}_{f}(x)=\left\{X\in T_{x}M:\alpha_{f}(X,Y)=0\;\;\text{for all}\;\,Y\in T_{x}M\right\}.$
The following result will be used in the sequel.
Proposition 10.
Let $f\colon M^{n}\to\mathbb{Q}_{c}^{n+m},n\geq 4,c\geq 0$, be a substantial isometric immersion of a compact, simply connected, even-dimensional manifold with flat normal bundle. Assume that equality holds in ($\ast$1) for $k=n/2$, and that there exist two distinct Dupin principal normals $\eta_{1}$ and $\eta_{2}$, each of multiplicity $k$, at every point $x\in M^{n}$ where $f$ is not umbilical, such that
$T_{x}M=E_{\eta_{1}}(x)\oplus E_{\eta_{2}}(x).$
Then, $M^{n}$ is isometric to a torus $\mathbb{S}^{k}(r)\times\mathbb{S}^{k}(\sqrt{R^{2}-r^{2}})$ and $f$ is a composition $f=j\circ g$, where $g\colon M^{n}\to\mathbb{S}^{n+1}(R)$ is the standard embedding of the torus $\mathbb{S}^{k}(r)\times\mathbb{S}^{k}(\sqrt{R^{2}-r^{2}})$ into a sphere $\mathbb{S}^{n+1}(R)$, and $j\colon\mathbb{S}^{n+1}(R)\to\mathbb{Q}_{c}^{n+2}$ is an umbilical inclusion, with $R\leq 1$ if $c=1$.
Proof.
Consider the open subset $M^{*}$ of points where $f$ is not umbilical. On $M^{*}$, the vector fields $\eta_{1}$ and $\eta_{2}$ are smooth Dupin principal normals, with smooth corresponding distributions $E_{1}=E_{\eta_{1}}$ and $E_{2}=E_{\eta_{2}}$, each of rank $k$ (see Proposition 1 in [21] or Lemma 10 in [28]).
Since
$2\mathcal{H}_{f}=\eta_{1}+\eta_{2}\;\;{\text{and}}\;\,S=k(\|\eta_{1}\|^{2}+\|\eta_{2}\|^{2}),$
condition ($\ast$1), which now holds as equality for $k=n/2$, implies
${\langle}\eta_{1},\eta_{2}{\rangle}=-c.$ (14)
The Codazzi equation for $f$ is easily seen to yield
${\langle}\nabla_{X}Y,Z{\rangle}(\eta_{i}-\eta_{j})={\langle}X,Y{\rangle}\nabla_{Z}^{\perp}\eta_{i}\;\,\text{if}\;\,i\neq j,$ (15)
for all $X,Y\in E_{i},Z\in E_{j},1\leq i,j\leq 2$. Using equation (14) and the fact that the Dupin principal normal vector fields $\eta_{1}$ and $\eta_{2}$ are parallel in the normal connection along $E_{1}$ and $E_{2}$, respectively (see Proposition 1.22 in [5]), it follows from equation (15) that the following relations hold on $M^{*}$:
${\langle}\nabla_{X}Y,Z{\rangle}(c+\|\eta_{j}\|^{2})=0\;\;\text{for all}\;\,X,Y\in E_{i}\;\;\text{and}\;\,Z\in E_{j},\;i\neq j.$ (16)
Now we distinguish two cases.
Case $c>0$. From equation (14), it follows that $M^{*}=M^{n}$. Consequently, equation (16) implies that the distributions $E_{1}$ and $E_{2}$ are totally geodesic on $M^{n}$. Since $M^{n}$ is simply connected, it is well-known that $M^{n}$ is a Riemannian product $M_{1}^{k}\times M_{2}^{k}$ (cf. Theorem $8.2$ in [5]) such that $TM_{i}^{k}=E_{i}$, $i=1,2$. As the second fundamental form of $f$ is adapted to the distributions $E_{1}$ and $E_{2}$, then Theorem $8.4$ and Corollary $8.6$ in [5] imply that the submanifold is an extrinsic product of isometric immersions, each of which is totally umbilical. Therefore, the submanifold is a torus $\mathbb{S}^{k}(r)\times\mathbb{S}^{k}(\sqrt{R^{2}-r^{2}})$ in a sphere $\mathbb{S}^{n+1}(R)\subset\mathbb{S}^{n+2}$.
Case $c=0$. Consider the open subset
$M_{+}=\left\{x\in M^{*}:\eta_{1}(x)\neq 0\,\;\text{and}\,\;\eta_{2}(x)\neq 0\right\}.$
Suppose that $M_{+}$ is nonempty. From (16), it follows that that the distributions $E_{1}$ and $E_{2}$ are totally geodesic on $M_{+}$. It is well-known that $M_{+}$ is locally a Riemannian product $M_{1}^{k}\times M_{2}^{k}$ (cf. Theorem $8.2$ in [5]) such that $TM_{i}^{k}=E_{i}$, $i=1,2$. Since the second fundamental form of $f$ is adapted to the distributions $E_{1}$ and $E_{2}$, Theorem $8.4$ and Corollary $8.6$ in [5] imply that the submanifold is locally an extrinsic product of isometric immersions, each of which is totally umbilical. Consequently, the submanifold is locally, on $M_{+}$, a torus $\mathbb{S}^{k}(r)\times\mathbb{S}^{k}(\sqrt{R^{2}-r^{2}})$ in a sphere $\mathbb{S}^{n+1}(R)\subset\mathbb{R}^{n+2}$.
Now, suppose that the interior $\operatorname{int}(M_{0})$ of the subset $M_{0}=M^{*}\smallsetminus M_{+}$ is nonempty. Assume that $\eta_{2}=0$ on a connected component $U$ of $\operatorname{int}(M_{0})$. Then $f$ has constant index of relative nullity $k$ on $U$. From (16), it follows that the relative nullity distribution $\mathcal{D}_{f}=E_{2}$ is totally geodesic. Let $C_{T}\colon\mathcal{D}_{f}^{\perp}\to\mathcal{D}_{f}^{\perp}$ be the associated splitting tensor for any $T\in\mathcal{D}_{f}$ (see [5, p. 186]). Since
$\alpha_{f}(X,Y)={\langle}X,Y{\rangle}\eta_{1}\;\;{\text{for all}}\;\,X\in\mathcal{X}(M^{n}),Y\in\mathcal{D}_{f}^{\perp},$
the Codazzi equation
[…]
implies that
$C_{T}={\langle}\nabla\log\|\eta_{1}\|,T{\rangle}Id,$
where $Id$ is the identity map on the conullity distribution $\mathcal{D}_{f}^{\perp}$. This shows that the conullity distribution is umbilical, and consequently, integrable.
Let $\varSigma^{k}$ be a leaf of $\mathcal{D}_{f}^{\perp}$. By Proposition 7.6 in [5], $f$ is locally a generalized cone over an isometric immersion $g\colon\varSigma^{k}\to\mathbb{Q}_{\tilde{c}}^{k+m}$ such that $f\circ j_{\varSigma}=i\circ g$, where $j_{\varSigma}\colon\varSigma^{k}\to M^{n}$ is the inclusion and $i\colon\mathbb{Q}_{\tilde{c}}^{k+m}\to\mathbb{R}^{n+m}$ is an umbilical inclusion. The second fundamental form of the immersion $f\circ j$ is given by
$\alpha_{f\circ j}(X,Y)={\langle}X,T{\rangle}\big{(}\eta_{1}+f_{*}\nabla\log\|\eta_{1}\|\big{)}\;\;{\text{for all}}\;\,X,Y\in\mathcal{X}(\varSigma^{k}).$
Thus, $f\circ j_{\varSigma}$ is umbilical and $g(\varSigma^{k})$ is a sphere $\mathbb{S}^{k}(r)$ centered at a point $x_{0}\in\mathbb{R}^{n+m}$ with radius
$r=\frac{1}{\sqrt{\|\eta_{1}\|^{2}+\|\nabla\log\|\eta_{1}\|\|^{2}}}.$
If the umbilical submanifold $\mathbb{Q}_{\tilde{c}}^{k+m}$ is totally geodesic in $\mathbb{R}^{n+m}$, then the submanifold $f$ is a $k$-cylinder over the sphere $\mathbb{S}^{k}(r)$. Alternatively, if $\mathbb{Q}_{\tilde{c}}^{k+m}$ is a sphere centered at a point $\tilde{x}_{0}\in\mathbb{R}^{n+m}$, then $f$ is a (k1)(k-1)-cylinder over a submanifold $N^{k+1}$, which is a cone over the sphere $\mathbb{S}^{k}(r)$ with its vertex at $\tilde{x}_{0}\neq x_{0}$.
Consequently, on $\operatorname{int}(M_{0})$ the submanifold is locally a $k$-cylinder in $\mathbb{R}^{n+1}$ over a sphere $\mathbb{S}^{k}(r)$, or a (k1)(k-1)-cylinder over a submanifold $N^{k+1}$ in $\mathbb{R}^{k+2}$, which is a cone over a sphere $\mathbb{S}^{k}(r)$.
In the later case, $f$ is given locally on $\operatorname{int}(M_{0})$ by
$f(x,w)=j(x)+w,\;\;(x,w)\in N_{i}\mathbb{S}^{k+m}(R),$
where the normal bundle $N_{i}\mathbb{S}^{k+m}(R)$ of the inclusion $i\colon\mathbb{S}^{k+m}(R)\to\mathbb{R}^{n+m}$ is regarded as subbundle of $N_{i\circ j}\mathbb{S}^{k}(r)$. Here, $j\colon\mathbb{S}^{k}(r)\to\mathbb{S}^{k+m}(R)$ is an umbilical inclusion with $r<R$. Equivalently, $f$ is locally parametrized by
$f(x,t_{0},t_{1},\dots,t_{k-1})=t_{0}j(x)+\sum_{i=1}^{k-1}t_{\ell}v_{\ell},\;\,t_{0}>0,\;x\in\mathbb{S}^{k}(r),$
where
$\mathbb{S}^{k+m}(R)\subset\mathbb{R}^{k+m+1},\;\;\mathbb{R}^{n+m}=\mathbb{R}^{k+m+1}\oplus\mathbb{R}^{k-1},$
and $\{v_{\ell}\}_{1\leq\ell\leq k-1}$ is an orthonormal basis of $\mathbb{R}^{k-1}$. The Laplacian operator $\Delta_{M}$ of $M^{n}$ is given by
$\Delta_{M}=\frac{k}{t_{0}}\frac{\partial}{\partial t_{0}}+\frac{\partial^{2}}{\partial t_{0}^{2}}+\frac{1}{t^{2}_{0}}\,\Delta_{\mathbb{S}^{k}(r)}+\sum_{\ell=1}^{k-1}\frac{\partial^{2}}{\partial t^{2}_{\ell}}.$
Since $S=a/t^{2}_{0}$, where $a$ is a positive constant, it follows that
$\Delta_{M}S=-\frac{a(n-6)}{t_{0}^{4}}.$
Consequently, either $\Delta_{M}S\geq 0$ or $\Delta_{M}S\leq 0$ on $\operatorname{int}(M_{0})$, depending on the dimension $n$.
Since $S$ is locally constant on $M_{+}$, continuity implies that $S$ is either superharmonic or subharmonic function on $M^{n}$. By the maximum principle, it follows that $S$ is a positive constant on $M^{n}$. Consequently, $M^{*}=M^{n}$, and therefore $M^{n}=M_{+}\cup M_{0}$.
Thus, on $M_{0}$, one of the Dupin principal normal vector fields vanishes, while the other has constant length $S/k$. On the other hand, both principal normal fields have constant and positive length on each connected component of $M_{+}$. By continuity and the compactness of $M^{n}$, it follows that $M_{+}=M^{n}$. Consequently, using a similar argument as in Case $c>0$, we conclude that $M^{n}$ is isometric to a torus $\mathbb{S}^{k}(r)\times\mathbb{S}^{k}(\sqrt{R^{2}-r^{2}})$, and the immersion $f$ is the standard embedding into a sphere $\mathbb{S}^{n+1}(R)\subset\mathbb{R}^{n+2}$. ∎
Proof of Theorem 2: Part $(i)$ of Proposition 8 implies that $M^{4}$ has nonnegative isotropic curvature and that the Bochner-Weitzenböck operator $\mathcal{B}^{[2]}$ is nonnegative at every point. The proof of the theorem is divided into three cases.
Case I. We begin by proving the theorem for simply connected submanifolds. Since $M^{4}$ has nonnegative isotropic curvature, Theorem 4.10 in [18] implies that one of the following holds:
  1. (a)
    $M^{4}$ carries a metric of positive isotropic curvature.
  2. (b)
    $M^{4}$ is diffeomorphic to a product $\mathbb{S}^{2}\times\varSigma^{2}$, where $\varSigma^{2}$ is a compact surface.
  3. (c)
    $M^{4}$ is a Kähler manifold biholomorphic to $\mathord{\mathbb{C}}\mathord{\mathbb{P}}^{2}$.
We analyse each case separately as follows.
Case (a). We assert that the manifold $M^{4}$ is diffeomorphic to $\mathbb{S}^{4}$. Given that $M^{4}$ is simply connected by assumption, the main result in [17] implies that $M^{4}$ is homeomorphic to $\mathbb{S}^{4}$. Furthermore, $M^{4}$ is locally irreducible. If this were not the case, then Theorem 3.1 in [18] would imply that $M^{4}$ is isometric to a Riemannian product of two compact surfaces, which leads to a contradiction. Hence, $M^{4}$ is locally irreducible, and by Theorem 2 in [3], one of the following cases must hold:
  1. (i)
    $M^{4}$ is diffeomorphic to a spherical space form.
  2. (ii)
    $M^{4}$ is a Kähler manifold biholomorphic to $\mathord{\mathbb{C}}\mathord{\mathbb{P}}^{2}$.
  3. (iii)
    $M^{4}$ is isometric to a compact symmetric space.
Since $M^{4}$ is homeomorphic to $\mathbb{S}^{4}$, case (ii) above is excluded. In case (i), the manifold is clearly diffeomorphic to $\mathbb{S}^{4}$. Moreover, the only compact symmetric spaces of dimension four are the round sphere, the product of two 2-dimensional spheres, the complex projective plane, and the quaternionic projective line $\mathbb{H}P^{1}\cong\mathbb{S}^{4}$. Consequently, in case (iii), the manifold $M^{4}$ must also be diffeomorphic to $\mathbb{S}^{4}$.
Case (b). Since $M^{4}$ is simply connected, the surface $\varSigma^{2}$ must be diffeomorphic to $\mathbb{S}^{2}$, which implies that $M^{4}$ is diffeomorphic to $\mathbb{S}^{2}\times\mathbb{S}^{2}$. Consequently, $\beta_{\pm}(M)=1$. It then follows from part $(ii)$ of Proposition 8 that equality holds in ($\ast$1). Moreover, at each point $x\in M^{4}$, there exists an oriented orthonormal four-frame $\{e_{1},e_{2},e_{3},e_{4}\}\subset T_{x}M$ such that the second fundamental form $\alpha_{f}$ of $f$ satisfies conditions (3)-(5).
In addition, there exist a nontrivial self-dual harmonic 2-form $\omega_{+}$ and a nontrivial anti-self-dual harmonic 2-form $\omega_{-}$. By the Bochner-Weitzenböck formula, both $\omega_{+}$ and $\omega_{-}$ are parallel and
${\langle}\mathcal{B}^{[2]}(\omega_{\pm}),\omega_{\pm}{\rangle}=0$
everywhere. Consequently, $\ker\mathcal{B}_{+}^{[2]}\neq 0$ and $\ker\mathcal{B}_{-}^{[2]}\neq 0$ at every point. Hence, both parts $(ii)$ and $(iii)$ of Lemma 9 hold everywhere.
We claim that the immersion $f$ has flat normal bundle at every point $x\in M^{4}$. Suppose first that both cases (ii1) and (iii1) hold simultaneously at $x$. Then, it follows from (3) and (5) that $c=0$, and hence $f$ is totally geodesic at this point. If, instead, either (ii1) or (iii1) does not hold at $x$, then Lemma 9 implies that $f$ has flat normal bundle at this point. This proves the claim that $f$ has flat normal bundle.
Since by part $(ii)$ of Proposition 8, the shape operator, associated with any normal direction, has at most two distinct eigenvalues, each of multiplicity $2$ at every point. This together with the flatness of the normal bundle implies that there exist two distinct Dupin principal normals, $\eta_{1}$ and $\eta_{2}$ each of multiplicity $2$, at every point $x\in M^{4}$ where $f$ is not umbilical, such that
$T_{x}M=E_{\eta_{1}}(x)\oplus E_{\eta_{2}}(x).$
It then follows from Proposition 10 that the submanifold is as described in part $(iii$-$a)$ of the theorem.
Case (c). In this case, either
$\beta_{+}(M)=1\;\;\text{and}\;\;\beta_{-}(M)=0,\;\;\text{or}\;\;\beta_{+}(M)=0\;\;\text{and}\;\;\beta_{-}(M)=1.$
We will treat only the former case, as the latter can be handled in a similar manner. It then follows from part $(ii)$ of Proposition 8 that equality holds in ($\ast$1). Moreover, at each point $x\in M^{4}$, there exists an oriented orthonormal four-frame $\{e^{x}_{1},e^{x}_{2},e^{x}_{3},e^{x}_{4}\}\subset T_{x}M$ such that the second fundamental form $\alpha_{f}$ of $f$ satisfies conditions (3)-(5).
Clearly, there exists a nontrivial self-dual harmonic 2-form $\omega_{+}$. By the Bochner-Weitzenböck formula, the form $\omega_{+}$ is parallel, and
${\langle}\mathcal{B}^{[2]}(\omega_{+}),\omega_{+}{\rangle}=0$
at every point. Hence, $\ker\mathcal{B}_{+}^{[2]}\neq 0$ at every point, and part $(ii)$ of Lemma 9 applies.
We consider the subset
$M_{1}=\left\{x\in M^{4}:\mu^{+}_{1}(x)=0\;\;\text{and}\;\;\mu^{+}_{2}(x)\mu^{+}_{3}(x)>0\right\},$
where $\mu^{+}_{1},\mu^{+}_{2}$, and $\mu^{+}_{3}$ are as given in Lemma 9. It follows that
$M_{1}=\left\{x\in M^{4}:\mu^{+}_{1}(x)=0\;\;\text{and}\;\;\|\alpha_{11}-\alpha_{44}\|^{2}+4\|\alpha_{13}\|^{2}+4\|\alpha_{14}\|^{2}>0\right\}.$
We claim that this subset is nonempty. Suppose, to the contrary, that $M_{1}$ is empty. Then, at every point of $M^{4}$, one of the following conditions must hold:
$\alpha_{14}+\alpha_{23}\neq 0,\;\text{or}\;\alpha_{13}-\alpha_{24}\neq 0,\;\text{or}\;\|\alpha_{11}-\alpha_{44}\|^{2}+4\|\alpha_{13}\|^{2}+4\|\alpha_{14}\|^{2}=0.$ (17)
We claim that the immersion $f$ has flat normal bundle. Indeed, if the third condition in (17) holds at a point $x$, then (3) and (4) imply that $c=0$, and hence $f$ is totally geodesic at $x$. If, instead, either the first or the second condition in (17) holds at a point $x$, then part $(ii)$ of Lemma 9 shows that either case (ii2) or case (iii2) holds at that point. In either situation, Lemma 9 further implies that the submanifold $f$ has flat normal bundle and satisfies
$\dim N_{1}f\leq 2$
at all points of $M^{4}$.
By part $(ii)$ of Proposition 8, the shape operator, associated with any normal direction, has at most two distinct eigenvalues, each of multiplicity $2$ at every point. Combined with the flatness of the normal bundle, this implies the existence of two distinct Dupin principal normals, $\eta_{1}$ and $\eta_{2}$, each of multiplicity $2$, at every point $x\in M$ where $f$ is not umbilical, such that
$T_{x}M=E_{\eta_{1}}(x)\oplus E_{\eta_{2}}(x).$
Since the submanifold is substantial, Proposition 10 implies that its codimension is $m=2$. Consequently, the complex projective plane $\mathord{\mathbb{C}}\mathord{\mathbb{P}}^{2}$ would admit an immersion into $\mathbb{R}^{6}$ or in $\mathbb{S}^{6}$. This, however, contradicts known topological obstructions (see, for instance, [7, pp. 290-291]). Hence, $M_{1}$ must be nonempty.
From part $(i)$ of Lemma 9, it follows that at every point of $M_{1}$ we have
$\mu_{1}^{+}=0,\;\;\mu_{2}^{+}=\mu_{3}^{+}>0.$
Consequently, the kernel of $\mathcal{B}_{+}^{[2]}$ at each point $x\in M_{1}$ is spanned by the vector $e^{x}_{1}\wedge e^{x}_{2}+e^{x}_{3}\wedge e^{x}_{4}$.
Let $Z$ be the dual to the self-dual form $\omega_{+}$. Since this form is parallel, we may normalize it (after multiplying by a constant if necessary) so that $\|\omega_{+}\|=\sqrt{2}$. Fix a point $x_{0}\in M_{1}$. At this point, we may assume that
$Z_{x_{0}}=e^{x_{0}}_{1}\wedge e^{x_{0}}_{2}+e^{x_{0}}_{3}\wedge e^{x_{0}}_{4}.$
Now, consider the almost complex structure $J_{x_{0}}\colon T_{x_{0}}M\to T_{x_{0}}M$ given by $J_{x_{0}}e^{x_{0}}_{1}=e^{x_{0}}_{2}$ and $J_{x_{0}}e^{x_{0}}_{3}=e^{x_{0}}_{4}$. Then, we have
$\omega_{+}(v,w)={\langle}J_{x_{0}}v,w{\rangle}\;\;\text{for all}\;\,v,w\in T_{x_{0}}M.$
Moreover, we define the skew-symmetric endomorphism $J$ of the tangent bundle of $M^{4}$ such that
$\omega_{+}(X,Y)={\langle}JX,Y{\rangle}\;\;\text{for all}\;\,X,Y\in\mathcal{X}(M).$
Clearly, $J$ is parallel because $\omega_{+}$ is parallel. We claim that $J$ is orthogonal; that is, $\|J_{x}v\|=\|v\|$ for every point $x\in M^{4}$ and every $v\in T_{x}M$. Indeed, let $V$ be a parallel vector field along a curve $c\colon[0,1]\to M$ such that $c(0)=x,c(1)=x_{0}$ and $V(0)=v$. Obviously, the vector field $W=JV$ is also parallel along $c$. Using that $J_{x_{0}}$ is orthogonal we have
$\|J_{x}v\|=\|W(0)\|=\|W(1)\|=\|J_{x_{0}}V(1)\|=\|V(1)\|=\|V(0)\|=\|v\|,$
which proves the claim. Since $J$ is both skew-symmetric and orthogonal, it defines an almost complex structure that is also parallel. Hence, the triple $(M^{4},{\langle}\cdot,\cdot{\rangle},J)$ is a Kähler manifold.
We now claim that, at every point $x\in M^{4}$, we have
$\alpha_{f}(J_{x}v,J_{x}w)=\alpha_{f}(v,w)\;\;\text{for all}\;\;v,w\in T_{x}M.$ (18)
To prove this, we distinguish two cases.
First, let $x$ be an arbitrary point in $M_{1}$. Since the kernel of $\mathcal{B}_{+}^{[2]}$ at $x$ is spanned by the vector $e^{x}_{1}\wedge e^{x}_{2}+e^{x}_{3}\wedge e^{x}_{4}$, and since $\|Z_{x}\|=\sqrt{2}$, it follows that
$Z_{x}=\pm\left(e^{x}_{1}\wedge e^{x}_{2}+e^{x}_{3}\wedge e^{x}_{4}\right).$
Because $Z$ is the dual of the self-dual form $\omega_{+}$, we have
${\langle}J_{x}v,w{\rangle}={\langle}{\langle}Z_{x},v\wedge w{\rangle}{\rangle}\;\;\text{for all}\;\,v,w\in T_{x}M.$
Consequently,
$J_{x}e^{x}_{1}=\pm e^{x}_{2}\;\;\text{and}\;\;J_{x}e^{x}_{3}=\pm e^{x}_{4}$
at each point $x\in M_{1}$. On the other hand, we have
$\alpha_{14}+\alpha_{23}=\alpha_{13}-\alpha_{24}=0$ (19)
at every point of $M_{1}$. Combining (3) with (19), we obtain
$\alpha_{f}(J_{x}e^{x}_{i},J_{x}e^{x}_{j})=\alpha_{ij}\;\;\text{for all}\;\;1\leq i,j\leq 4,$ (20)
and thus (18) follows from (20) by linearity.
Having analyzed the case $x\in M_{1}$, we now turn to the complementary case $x\notin M_{1}$. In this case, (17) holds at $x$.
If the third condition in (17) is satisfied, then (3) and (5) imply that $c=0$, and hence $f$ is totally geodesic at $x$. In this situation, (18) holds trivially.
Next, suppose that the first or the second condition in (17) holds at $x$. We treat only the first case, since the second can be handled in a similar way, and will therefore be omitted. Then part $(ii)$ of Lemma 9 shows that case (ii2) holds at that point. Together with (ii2), it follows from part $(i)$ of this lemma that the matrix of the Bochner-Weitzenböck operator $\mathcal{B}_{+}^{[2]}$ at $x$, with respect to the basis $\{\eta_{i}\}_{1\leq i\leq 3}$ of $\Lambda_{+}^{2}T_{x}M$, is given by
$4\|\alpha_{14}\|^{2}\begin{bmatrix}1&\rho&0\\ \rho&\rho^{2}&0\\ 0&0&\rho^{2}+1&\!\!\!\!\!\end{bmatrix}$,
where $\rho$ is the constant appearing in case (ii2) of Lemma 9. Therefore, the kernel of $\mathcal{B}_{+}^{[2]}$ at $x$ is spanned by the vector $\rho\eta_{1}+\eta_{2}$, and since $\|Z_{x}\|=\sqrt{2}$, it follows that
$Z_{x}=\lambda\left(-\rho\eta_{1}+\eta_{2}\right),$
where the number $\lambda$ satisfies
$\lambda^{2}=\frac{2}{\rho^{2}+1}.$ (21)
Because $Z$ is the dual of the self-dual form $\omega_{+}$, we have
${\langle}J_{x}v,w{\rangle}={\langle}{\langle}Z_{x},v\wedge w{\rangle}{\rangle}\;\;\text{for all}\;\,v,w\in T_{x}M.$
Using this relation together with (21), we obtain
$J_{x}e^{x}_{1}$ $=\frac{\lambda}{\sqrt{2}}\left(-\rho\,e^{x}_{2}+e^{x}_{3}\right),\;\;J_{x}e^{x}_{2}=\frac{\lambda}{\sqrt{2}}\left(\rho\,e^{x}_{1}-e^{x}_{4}\right),$
$J_{x}e^{x}_{3}$ $=-\frac{\lambda}{\sqrt{2}}\left(e^{x}_{1}+\rho\,e^{x}_{4}\right),\;\;J_{x}e^{x}_{4}=\frac{\lambda}{\sqrt{2}}\left(e^{x}_{2}+\rho\,e^{x}_{3}\right).$
From these expressions, and combining (3) with (ii2), we obtain (20). Hence (18) follows directly.
Therefore, the second fundamental form of the submanifold satisfies
$\alpha_{f}(JX,JY)=\alpha_{f}(X,Y)\;\;\text{for all}\;\,X,Y\in\mathcal{X}(M).$ (22)
Immersions satisfying condition (22) have parallel second fundamental form (see Theorem 4 in [9]). Moreover, under the immersion $f$, each geodesic of $M^{4}$ is mapped into a plane circle. Such submanifolds were classified in [9] and [22]. From this classification, it follows that the submanifold is as described in part $(iii$-$b)$ of the theorem, thereby completing the proof for the case where the manifold is simply connected.
Case II. Now suppose that the fundamental group of $M^{4}$ is finite. Then we claim that $M^{4}$ is simply connected. Consider the universal covering $\pi\colon\tilde{M}^{4}\to M^{4}$. Since the fundamental group of $M^{4}$ is finite, $\tilde{M}^{4}$ must be compact. Moreover, the isometric immersion $\tilde{f}=f\circ\pi$ satisfies ($\ast$1). Therefore, from Case I, we conclude that either $\tilde{M}^{4}$ is diffeomorphic to $\mathbb{S}^{4}$, or the submanifold $\tilde{f}$ is as described in parts $(i)$ or $(ii)$ of the theorem.
Assume first that $\tilde{M}^{4}$ is diffeomorphic to $\mathbb{S}^{4}$. Arguing as in the proof of Case I, we conclude that $M^{4}$ is locally irreducible. Then, by Theorem 2 in [3], $M^{4}$ is either diffeomorphic to a spherical space form or isometric to a compact symmetric space. In the former case, it is clear that the manifold $M^{4}$ must be diffeomorphic to $\mathbb{S}^{4}$. In the latter case, since the only 4-dimensional compact symmetric spaces are the round spheres, the product of two 2-dimensional spheres, the complex plane, and the quaternionic projective line $\mathbb{H}P^{1}$, it follows that $M^{4}$ must also be diffeomorphic to $\mathbb{S}^{4}$.
Now, suppose that the submanifold $\tilde{f}$ is as described in parts $(i)$ or $(ii)$ of the theorem. In either case, the covering map $\pi\colon\tilde{M}^{4}\to M^{4}$ must be a diffeomorphism. This observation completes the proof of the theorem in this case.
Case III. Finally, suppose that the fundamental group $\pi_{1}(M)$ of $M^{4}$ is infinite. We first claim that $M^{4}$ is locally reducible. If this were not the case, then by Theorem 2 in [3], it would follow that either $M^{4}$ is diffeomorphic to a spherical space form, or its universal cover $\tilde{M}^{4}$ is biholomorphic to $\mathord{\mathbb{C}}\mathord{\mathbb{P}}^{2}$, or isometric to a compact symmetric space. Each of these possibilities contradicts the fact that $\tilde{M}^{4}$ is not compact. Therefore, $M^{4}$ must be locally reducible. It then follows from Theorem 3.1 in [18] that the universal cover $\tilde{M}^{4}$ is isometric to one of the following Riemannian products:
  1. (i)
    $(\mathbb{R}^{n_{0}},g_{0})\times(N_{1}^{n_{1}},g_{1})\times(N_{2}^{n_{2}},g_{2})$, where $n_{0}\geq 0,g_{0}$ is the flat Euclidean metric, and either $n_{i}=2$ and $N_{i}=\mathbb{S}^{2},i=1,2$, has nonnegative Gaussian curvature, or else $n_{i}=3$ and $N_{i}$ is compact with nonnegative Ricci curvature.
  2. (ii)
    $(\varSigma^{2},g_{\varSigma})\times(N^{2},g_{N})$, where $\varSigma^{2}$ is a surface whose Gaussian curvature $K_{\varSigma}$ is negative at some point, and $N^{2}$ is a compact surface with positive Gaussian curvature $K_{N}$.
First, we claim that case (ii) above cannot occur. Suppose, for the sake of contradiction, that it does. Clearly the immersion $\tilde{f}=f\circ\pi$ satisfies ($\ast$1), where $\pi\colon\tilde{M}^{4}\to M^{4}$ is the covering map. We consider the isometric immersions
$f_{\varSigma}=\tilde{f}\circ\mathfrak{i}_{\varSigma}\;\;\text{and}\;\;f_{N}=\tilde{f}\circ\mathfrak{i}_{N},$
where $\mathfrak{i}_{\varSigma}\colon\varSigma\to\tilde{M}$ and $\mathfrak{i}_{N}\colon N\to\tilde{M}$ are the inclusion maps. Since both immersions $\mathfrak{i}_{\varSigma},\mathfrak{i}_{N}$ are totally geodesic, it follows that the second fundamental form of $f_{\varSigma}$ and $f_{N}$ are given, respectively, by
$\alpha_{f_{\varSigma}}(X,Y)$ $=\alpha_{\tilde{f}}(\mathfrak{i}_{\varSigma_{*}}X,\mathfrak{i}_{\varSigma_{*}}Y),\;X,Y\in T\varSigma,$
$\alpha_{f_{N}}(V,W)$ $=\alpha_{\tilde{f}}(\mathfrak{i}_{N_{*}}V,\mathfrak{i}_{N_{*}}W),\;V,W\in TN.$
Given that $\tilde{M}^{4}$ is isometric to the Riemannian product $(\varSigma^{2},g_{\varSigma})\times(N^{2},g_{N})$, it follows from the above that
$S_{\tilde{f}}$ $=S_{f_{\varSigma}}+S_{f_{N}}+2\sum_{i=1,2}\sum_{j=3,4}\|\alpha_{\tilde{f}}(\mathfrak{i}_{\varSigma_{*}}e_{i},\mathfrak{i}_{N_{*}}e_{j})\|^{2}$ (23)
and
$4\tilde{H}^{2}$ $=H^{2}_{f_{\varSigma}}+H^{2}_{f_{N}}+2{\langle}\mathcal{H}_{f_{\varSigma}},\mathcal{H}_{f_{N}}{\rangle},$ (24)
where $\{e_{i}\}_{1\leq i\leq 4}$ is an orthonormal frame such that $e_{1},e_{2}\in T\varSigma$, $e_{3},e_{4}\in TN$, and $\tilde{H}$ is the mean curvature of the immersion $\tilde{f}$. The squared lengths of the traceless parts $\Phi_{f_{\varSigma}}$ and $\Phi_{f_{N}}$ of the second fundamental forms of the immersions $f_{\varSigma}$ and $f_{N}$ are given by
$\|\Phi_{f_{\varSigma}}\|^{2}=S_{f_{\varSigma}}-2H^{2}_{f_{\varSigma}}\;\;{\text{and}}\;\;\|\Phi_{f_{N}}\|^{2}=S_{f_{N}}-2H^{2}_{f_{N}},$
respectively. Taking into account (23), (24) and the above, inequality ($\ast$1) for $\tilde{f}$ can be equivalently written as
$\|\Phi_{f_{\varSigma}}\|^{2}+\|\Phi_{f_{N}}\|^{2}+2\sum_{i=1,2}\sum_{j=3,4}\|\alpha_{\tilde{f}}(\mathfrak{i}_{\varSigma_{*}}e_{i},\mathfrak{i}_{N_{*}}e_{j})\|^{2}\leq 4\big{(}c+{\langle}\mathcal{H}_{f_{\varSigma}},\mathcal{H}_{f_{N}}{\rangle}\big{)}.$ (25)
On the other hand, from the Gauss equation and ($\ast$1), it follows that the scalar curvature $\tau$ of $\tilde{M}^{4}$ satisfies
$\tau\geq 8(c+\tilde{H}^{2}).$
Since $\tau=2(K_{\varSigma}+K_{N})$, we then obtain the inequality
$K_{\varSigma}+K_{N}\geq 4(c+\tilde{H}^{2}).$ (26)
Using (24) and the Gauss equation for both $f_{\varSigma}$ and $f_{N}$, we find that (26) is written as
$\|\Phi_{f_{\varSigma}}\|^{2}+\|\Phi_{f_{N}}\|^{2}+4\big{(}c+{\langle}\mathcal{H}_{f_{\varSigma}},\mathcal{H}_{f_{N}}{\rangle}\big{)}\leq 0.$ (27)
Then, from (25) and (27) we obtain $\Phi_{f_{\varSigma}}=0$. Hence, the immersion $f_{\varSigma}$ it totally umbilical, and consequently $K_{\varSigma}\geq 0$, which is clearly a contradiction.
Thus, only case (i) holds. Clearly, $n_{0}\neq 3$. We now claim that $n_{0}=1$. If $n_{0}=4$, then $M^{4}$ is flat. By the Gauss equation, we obtain
$S=12c+16H^{2}.$
Together with ($\ast$1), this implies that $c=0$ and $f$ is minimal, which leads to a contradiction. If $n_{0}=0$ or $n_{0}=2$, then the curvature operator of $M^{4}$ is nonnegative. In this case, by Theorem 1.3 in [10], $M^{4}$ is diffeomorphic to a quotient of one of the spaces $\mathbb{S}^{4}$, $\mathord{\mathbb{C}}\mathord{\mathbb{P}}^{4}$, $\mathbb{R}^{1}\times\mathbb{S}^{3}$, $\mathbb{S}^{2}\times\mathbb{S}^{2}$, $\mathbb{R}^{2}\times\mathbb{S}^{2}$, or $\mathbb{R}^{4}$ by a finite group of fixed-point-free isometries in the standard metric. This contradicts the fact that $\pi_{1}(M)$ is infinite. Therefore, $n_{0}=1$ and $\tilde{M}$ splits isometrically as $\mathbb{R}\times N^{3}$, where $N^{3}$ is compact, simply connected, and has nonnegative Ricci curvature. From Theorem 1.2 in [10], we then conclude that $N^{3}$ is diffeomorphic to $\mathbb{S}^{3}$.∎
Proof of Theorem 3: Let $n\geq 5$, and suppose that the submanifold $f$ is not as described in parts (i)(i)-$(iii)$ of Theorem 1. Then, the homology groups of $M^{n}$ satisfy
$H_{p}(M^{n};\mathbb{Z})=0\;\,\text{for all}\;\,2\leq p\leq n-2\;\;\text{and}\;\,H_{1}(M^{n};\mathbb{Z})\cong\mathbb{Z}^{\beta_{1}(M)}.$
Since $M^{n}$ is oriented, we have $H^{n}(M^{n};\mathbb{Z})\cong\mathbb{Z}$. Applying the universal coefficient theorem, we deduce that $H_{n-1}(M^{n};\mathbb{Z})$ is torsion-free. By Poincaré duality, it follows that $H_{n-1}(M^{n};\mathbb{Z})\cong\mathbb{Z}^{\beta_{1}(M)}$.
In the case $n=4$, the result follows directly from Theorem 2.∎
3.1 Examples of submanifolds satisfying condition ($\ast$1)
We now describe a method for constructing geometrically distinct submanifolds that are diffeomorphic either to the sphere $\mathbb{S}^{n}$ or to the torus $\mathbb{S}^{1}\times\mathbb{S}^{n-1}$, and which, in addition, satisfy condition ($\ast$1) in the case $c=0$.
Proposition 11.
Let $g\colon N^{n-1}\to\mathbb{R}^{m_{1}},n\geq 4$, be an isometric immersion of a manifold $N^{n-1}$ satisfying condition ($\ast$1) for an integer $k=\ell-1$, with $2\leq\ell\leq n-2$, as a strict inequality at every point; that is,
$S_{g}<a(n-1,\ell-1,H_{g},0).$
Consider any closed unit-speed curve $\gamma\colon\mathbb{S}^{1}\to\mathbb{R}^{m_{2}}$ whose first curvature $\kappa_{1}$ satisfies
$\kappa_{1}^{2}\leq\frac{n-\ell}{n-\ell-1}\Big{(}\min\big{(}a(n-1,\ell-1,H_{g},0)-S_{g}\big{)}\Big{)}.$ (28)
Then, the product immersion
$f=\gamma\times g\colon\mathbb{S}^{1}\times N^{n-1}\to\mathbb{R}^{m_{1}+m_{2}}$
satisfies condition ($\ast$1) for $k=\ell$.
Proof.
The squared length $S_{f}$ of the second fundamental form, and the mean curvature $H_{f}$ of the product immersion $f$ are given by
$S_{f}=\kappa_{1}^{2}+S_{g}\;\;{\text{and}}\;\,n^{2}H^{2}_{f}=\kappa_{1}^{2}+(n-1)^{2}H^{2}_{g}.$
From these expressions, it follows that condition ($\ast$1) for $f$, with $k=\ell$ and $c=0$, is equivalent to inequality (28), thereby completing the proof.∎
Next, we show that there exist many isometric immersions $g\colon N^{n-1}\to\mathbb{R}^{m}$ satisfying the hypotheses of Proposition 11. In particular, we construct geometrically distinct immersions of $\mathbb{S}^{n}$ into the Euclidean space $\mathbb{R}^{n+1}$ that satisfy condition ($\ast$1) for all $n\geq 3$ and $1\leq k\leq n-1$, either with strict inequality or with equality.
Proposition 12.
Let $f\colon M^{n}\to\mathbb{R}^{n+1},n\geq 3$, be an ovaloid in the Euclidean space with principal curvatures $0<\lambda_{1}\leq\dots\leq\lambda_{n}$. If
$\max\lambda_{n}\leq\min\lambda_{1}\Big{(}{\frac{n}{n-k}}\Big{)}^{1/2}$ (29)
for some integer $1\leq k\leq n-1$, then $f$ satisfies condition ($\ast$1) at every point. Moreover, if inequality (29) is strict, then inequality ($\ast$1) is also strict at every point.
Proof.
Since
$a(n,k,H,0)=\frac{n^{2}H^{2}}{n-k},$
the proof follows directly from the inequalities $S\leq n\lambda^{2}_{n}$ and $H\geq\lambda_{1}$.∎
A large class of ellipsoids satisfies condition (29). Consider, for instance, the ellipsoid in $\mathbb{R}^{n+1}$ defined by
$\frac{x_{1}^{2}}{a_{1}^{2}}+\cdots+\frac{x_{n+1}^{2}}{a_{n+1}^{2}}=1,$
where $0<a_{1}\leq\dots\leq a_{n+1}$. A straightforward computation shows that the minimum and the maximum of the principal curvatures of the ellipsoid are $a_{1}/a_{n+1}^{2}$ and $a_{n+1}/a_{1}^{2}$, respectively. It follows that condition (29) holds provided
$a_{n+1}\leq a_{1}\big{(}{\frac{n}{n-k}}\big{)}^{1/6}.$
By Hadamard’s classical theorem, every ovaloid in $\mathbb{R}^{n+1}$ is diffeomorphic to $\mathbb{S}^{n}$. Consequently, Propositions 11 and 12 provide geometrically distinct isometric immersions of manifolds diffeomorphic either to the sphere $\mathbb{S}^{n}$ or to the torus $\mathbb{S}^{1}\times\mathbb{S}^{n-1}$ for $n\geq 4$, which also satisfy condition ($\ast$1) with $c=0$ and $k=2$. Thus, we obtain numerous compact, geometrically distinct submanifolds that strictly satisfy ($\ast$1) at every point. Moreover, this strict form is preserved under sufficiently small smooth deformations of any such example. Hence our results are optimal.
Example 13.
Consider a rotational hypersurface in $\mathbb{R}^{n+1},n\geq 4$, obtained by rotating a curve $c$ in the $x_{1}x_{2}$-plane around the $x_{1}$-axis. Suppose the curve is given as a graph of a positive function $u=u(x_{1})$. After choosing the orientation appropriately, the rotational hypersurface has two principal curvatures given by
$\lambda=\frac{1}{u(1+(u^{\prime})^{2})^{1/2}},\;\,\mu=-\frac{u^{\prime\prime}}{(1+(u^{\prime})^{2})^{3/2}},$
where $\lambda$ has multiplicity n1n-1. The hypersurface satisfies condition ($\ast$1) with $k=2$ in strict form provided that
$(n-1)\lambda^{2}+2(n-1)\lambda\mu-(n-3)\mu^{2}>0,$
or equivalently,
$a_{n}(1+(u^{\prime})^{2})<uu^{\prime\prime}<b_{n}(1+(u^{\prime})^{2}),$
where
$a_{n}=\frac{\sqrt{2(n-1)(n-2)}+n-1}{n-3},\;\,b_{n}=\frac{\sqrt{2(n-1)(n-2)}-(n-1)}{n-3}.$
By connecting such hypersurfaces to spheres, as in Example 3.3 in [16], we obtain compact hypersurfaces that satisfy condition ($\ast$1) with $k=2$, whose first Betti number can be any nonnegative integer. The homology of these hypersurfaces is precisely as described in part (iii) of Theorem 3.